### v1

# Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models

Abstract
As Large Language Models (LLMs) continue to grow in size, their computational and memory requirements increase correspondingly. Consequently, the exploration of cost-effective and efficient fine-tuning methods has become increasingly important. Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs. In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA. However, these methods also make the fine-tuning initialization process more complex, and it remains challenging to surpass the performance ceiling of full fine-tuning. To address these issues, this paper introduces an innovative method called Bone (Block Affine), which not only reduces memory overhead but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting. Experimental comparisons across two different LLM architectures (LLaMA2, RWKV6) and various parameter scales demonstrate that the Bone structure can achieve rapid convergence and superior data fitting without the need for complex initialization. For example, when fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8k and math benchmarks, Bone achieved fine-tuning scores of 49.36 and 8.8, respectively, outperforming PISSA by 5.84% and 1.96%.
1 Introduction
Large models have been integrated into various industries, revolutionizing many traditional technologies Radford et al. (2019); Raffel et al. (2020). However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios. Full-scale fine-tuning of large models is extremely costly; for example, training the LLaMA2-7B Touvron et al. (2023) model with bfloat16 Wang & Kanwar (2019) precision requires around 60GB of VRAM. In contrast, by using PEFT (Parameter-Efficient Fine-Tuning)Xu et al. (2023) techniques, fine-tuning can be done with just a single 24GB VRAM GPU. As a result, numerous PEFT techniques and quantization methods have emerged to reduce the training costs of large models. LoRA (Low-Rank Adaptation) Hu et al. (2021) has become one of the most popular PEFT methods due to its small parameter size, its ability to optimize the model’s main weights, and the fact that it doesn’t add extra inference overhead.
LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$. Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The Figure 2(a) illustrates the structure visualization. However, extensive experiments Ding et al. (2023); Liu et al. (2024b); Biderman et al. (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning. This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$. To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al. (2024), PISSA Meng et al. (2024), and LoRA-GAWang et al. (2024). Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility. For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants and surpass full fine-tuning. It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights. To calculate the information interactions within the weight matrix $W$, Bone employs an initially zeroed, trainable matrix, named bone, to perform Block Affine operations on $W$. The specific calculations of Block Affine are detailed in Section 3, and the structural visualization is shown in Figure  2. To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al. (2023), RWKV6 Peng et al. (2024)). As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  1, Table  2).
Advantages of Bone::
1. Bone does not require the complex initialization needed by LoRA variants, yet it achieves fast convergence and better data fitting.
2. Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
3. Bone provides a new direction for the development of PEFT techniques.
(a) llama2-7b
(b) RWKV6
Figure 1: The left image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation. The right image displays the loss curves for RWKV6 with different parameter settings, also fine-tuned on MetaMathQA. Both models were trained with a batch size of 64 for a full epoch. Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings. Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
2 Related Works
The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al. (2019), prefix tuningLiu et al. (2023), prompt tuningBrown (2020), LoRAHu et al. (2021), and layer-freezing methods such as LISA.
The adapter method does not require fine-tuning all the parameters of the pre-trained model. Instead, it introduces a small number of task-specific parameters to store knowledge related to that task, thereby reducing the computational demands of model fine-tuning. Prefix tuning is a lightweight fine-tuning method for generative tasks. It adds a continuous, task-specific vector sequence, called a prefix, to the input. Unlike prompts, prefixes are entirely composed of free parameters and do not correspond to actual tokens. Compared to traditional fine-tuning, prefix tuning only optimizes the prefix, meaning we only need to store a large Transformer model and a small set of task-specific prefixes, resulting in minimal overhead for additional tasks.
Prompt tuning defines a unique prompt for each task, appending it to the input data while freezing the pre-trained model during training. LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers, while keeping the pre-trained weights unchanged. This approach significantly addresses the inference latency introduced by adapters and the fitting challenges of prefix tuning.
In LoRA, the adapter matrices $A$ and $B$ are updated with the same learning rate, but using the same rate for both may not effectively learn the features. LoRA+ corrects this suboptimality by setting different learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency. The DoRA Liu et al. (2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency. PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction. Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning. OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal. This orthogonal initialization helps maintain the stability of the parameter space during optimization. LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
3 Bone: Block Affine
In this section, we will introduce the Bone method. Unlike the LoRA series of architectures, Bone employs a single learnable matrix, initialized to zero and named ”bone.” This matrix interacts with the model’s matrix $W$ through Block Affine operations to represent $\Delta W$. This approach ensures that the initial state remains unchanged while also facilitating the integration of information between the weights. Assuming the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$. The specific calculation for Block Affine is as follows:
$Y=X(W+\Delta W)$ (1)
$\Delta W_{n,n}=W_{n,n}bone_{n,n}+bone_{n,n}$ (2)
Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices. We conducted ablation experiments to verify the effectiveness of using a single learnable matrix structure. The comparison formulas are as follows:
$\Delta W_{n,n}=W_{n,n}bone\_A_{n,n}+bone\_B_{n,n}$ (3)
We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns are not equal. Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principles of fine-tuning. Therefore, we adjust $bone$ to be a low-rank matrix. When the weight matrix $W$ has dimensions $(n,m)$, it is divided according to a block_size $b$ into $W(n/b,b,m/b,b)$, yielding $(n/b)\times(m/b)$ blocks. Inspired by GQA(Grouped Query Attention)Ainslie et al. (2023), we group these blocks together to significantly reduce the number of trainable parameters without affecting performance. Different groups share a single Bone matrix for computation. Thus, the Bone matrix is set to $(g,b,b)$, where $g$ is the group size. The $W$ matrix is reshaped to $(gn,g,b,b)$, where $gn$ represents the number of groups. The Bone structure is illustrated in Figure 2(b). The final formula can be expressed as:
$W_{gn,g,b,b}=Reshape(W_{n,m})$ (4)
$\Delta W_{n,m}=Reshape(W_{gn,g,b,b}bone_{g,b,b}+bone_{g,b,b})$ (5)
During our comparative experiments between RWKV6 and LLaMA2, we found that the matrix $W$ has several different dimensions. For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$. These differences require us to continuously adjust g to ensure the structure functions correctly. To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, we modified the dimensions of Bone to $(n,b)$, where $b$ remains the size of the smallest block, and $n$ corresponds to the $n$ dimension of $W$. Thus, $g=n/b$. During computation, Bone only requires a dimension transformation, The Bone structure is illustrated in Figure 2(c), and the modified formula is as follows:
$W_{m/b,n/b,b,b}=Reshape(W_{n,m})$ (6)
$bone_{n/b,b,b}=Reshape(bone_{n,b})$ (7)
$\Delta W_{n,m}=Reshape(W_{m/b,n/b,b,b}bone_{n/b,b,b}+bone_{n/b,b,b})$ (8)
In this way, we can easily adapt to any dimension of W and models with different architectures. The compatibility method used above groups by columns, and naturally, we can also group by rows. The visualization is shown in Figure  2(d), and the formula is as follows:
$W_{n/b,m/b,b,b}=Reshape(W_{n,m})$ (9)
$bone_{m/b,b,b}=Reshape(bone_{m,b})$ (10)
$\Delta W_{n,m}=Reshape(W_{n/b,m/b,b,b}bone_{m/b,b,b}+bone_{m/b,b,b})$ (11)
(a) LoRA
(b) Bone-both
(c) Bone-col
(d) Bone-row
[…]
4 Experiments
The experiments were conducted on 4×NVIDIA 4090 24G GPUs.In our experiments, we fixed lora_r=32 to ensure consistent trainable parameters across different fine-tuning methods, using the AdamW optimizer with a batch size of 64, a learning rate of 2e-5 and a warmup ratio of 0.0, without any weight decay.
4.1 Evaluating the Performance of Bone under Different Parameter Settings in RWKV6
Table  1presents a comparative evaluation of the fine-tuned RWKV6 models with different parameter settings on the NLG task. I fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al. (2023) and evaluated them on GSM8k Cobbe et al. (2021) and MATH datasets Yu et al. (2023). To ensure the accuracy of the experiment, the comparison between LoRA, PiSSA, and Bone was conducted with the same number of trainable parameters. The fine-tuning experiments for the 1.6B, 3B, and 7B RWKV6 models were each conducted for a complete epoch with a total batch size of 64 and a fixed learning rate of 2e-5.
Table 1: The performance of different PEFT techniques under various parameter settings on RWKV6.
Model Strategy Trainable GSM8K Math
RWKV 6-1.6B LoRA 33.0M 10.84 1.14
PiSSA 33.0M 12.89 1.28
Bone 33.0M 14.71 1.94
RWKV 6-3B LoRA 55.1M 21.22 2.34
PiSSA 55.1M 24.64 2.54
Bone 55.1M 25.25 2.82
RWKV 6-7B LoRA 88.1M 38.13 6.06
PiSSA 88.1M 40.48 6.12
Bone 88.1M 42.76 6.34
As shown in Table 1, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning. For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes. This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
4.2 Experiments on Llama2-7B
Models and Datasets
[…]
1. Math: We train our model on a 395k subset of MetaMathQA Yu et al. (2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al. (2021) and MATH  Yu et al. (2023), with higher complexity and diversity.
2. Code: We train our model on a 100k subset of CodeFeedback Zheng et al. (2024b), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval  Chen et al. (2021).
3. Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct Xu et al. (2024). We test our model on the MT-Bench dataset  Zheng et al. (2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects. The quality of the responses is judged by GPT-4o, and we report the first turn score.
Result
The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other structures in terms of convergence speed, data fitting, and generalization capabilities. Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further demonstrating the generalizability of the Bone approach.
Table 2: We fine-tuned LLaMA2-7B using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
Model Strategy Trainable GSM8K Math HumanEval MT-Bench
Llama2-7B LoRA 89.9M 38.97 0.44 16.46 3.93
OLoRA 89.9M 41.93 6.48 20.12 4.09
PiSSA 89.9M 43.52 6.92 20.73 4.25
Bone 87.0M 49.36 8.88 25.62 4.81
4.3 Experiments were conducted with different values of block_size $(b)$.
This subsection explores the upper limits of the Bone structure by varying the block_size in the Bone matrix. Notably, while we change the size of $b$, the value of $n$ remains constant. Therefore, as $b$ increases, the dimensions of Bone transform into the bone_g form, leading to a corresponding reduction in the number of blocks within each group. Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  3, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when b reaches 16, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is weaker, indicating that the value of b plays a crucial role in determining the model’s generalization ability. As b increases, the model’s generalization capacity strengthens.
Table 3: Comparing different values of block_size $(b)$
Model Strategy Trainable GSM8K Math
Llama2-7B LoRA 89.9M 38.97 0.44
PiSSA 89.9M 43.52 6.92
Bone b=16 21.7M 45.79 3.54
Bone b=32 43.5M 46.70 7.67
Bone b=64 87.0M 49.36 8.88
Bone b=128 174.0M 53.15 10.56
Figure 3: Training loss curves of Bone with different block_size on the MetaMathQA dataset.
4.4 Ablation experiments on different grouping methods.
In this subsection, we explore the impact of different grouping methods in the Bone structure on model fine-tuning performance. Due to structural differences in the weight matrix W, the Bone-free grouping requires manual configuration, which is inconvenient. Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure. We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math. The results are shown in Table  4. Since in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bone structure, making it difficult to align parameter counts. Even though Bone-row performs better than Pissa with only around 20M parameters, the gap between Bone-row and Bone-col is too large. Initially, we suspected this gap was due to Bone-row lacking about 14M parameters compared to Bone-col. However, this assumption was disproved when we observed that Bone-row with 32 blocks and 43.5M parameters performed similarly to Bone-row with 72.8M parameters. This made it difficult to directly evaluate which grouping method performs better.
To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bone fine-tuning. This allows for a fairer comparison between Bone-row and Bone-col. The experimental results, shown in Table  4, indicate that the difference between the two is minimal, with both performing well. Therefore, we believe that Bone can effectively fit data regardless of the grouping method used. The key factor influencing Bone’s performance remains the block size. However, this doesn’t imply that different grouping methods are meaningless. As more LLMs begin to use techniques like GQA and MLA Liu et al. (2024a) to reduce KV cache overhead Dai et al. (2024); Lee et al. (2024); Shazeer (2019), the main weight matrices become smaller, and Bone will need to adjust its grouping or employ other techniques to adapt to these new technologies.
Table 4: Comparing Bone-row, Bone-col on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone-row 72.8M 45.76 7.82
Bone-col 87.0M 49.36 8.88
RWKV6-3B Bone-row 44.1M 25.93 3.32
Bone-col 44.1M 25.25 2.82
4.5 Ablation experiments on Block Affine.
In this subsection, we aim to verify the effectiveness of a single learnable matrix, and we name the comparison structure Bone-AB. It consists of two low-rank matrices and is grouped by rows. To ensure the trainable parameter count is consistent, we set $b=32$ in Bone-AB and $b=64$ in Bone-row. For this comparison experiment, we continue to evaluate the Math task, with the results shown in Table  5. Bone-AB performed the worst, further demonstrating Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
Table 5: Block Affine on math tasks
Model Strategy Trainable GSM8K Math
Llama2-7B Bone-col b=32 43.5M 46.70 7.67
Bone-row b=64 72.8M 45.76 7.82
Bone-AB b=32 72.8M 44.35 6.16
5 Resource and efficiency
Table  6 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU. The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512. Due to the structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage by 5GB compared to LoRA. Additionally, Bone’s token throughput is only two-thirds of LoRA’s. To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput.
Despite the fact that the Bone structure significantly outperforms LoRA variants in fine-tuning, we cannot ignore the efficiency issues. Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
Table 6: Resource and efficiency
Model Strategy Trainable GPU Memory Token throughput
RWKV-1.6B LoRA 33.0M 7314 MB 5.86 kt/s
Bone 33.0M 12276 MB 4.11 kt/s
Bone(checkpoint) 33.0M 7398 MB 3.46 kt/s
RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
Bone 55.1M 22978 MB 2.16 kt/s
Bone(checkpoint) 55.1M 12168 MB 1.72 kt/s
6 Conclusion
This paper introduces a new PEFT technique called Bone(Block Affine), which employs block affine transformations on the weight matrix to capture the internal correlations within $W$, while also avoiding the complexity of initialization. Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA. Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility. Bone brings new possibilities to existing LLM PEFT techniques. Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
7 Limitations and future work
Due to limited training resources, this paper lacks comparative experiments on full fine-tuning. However, the extensive experiments presented demonstrate that Bone possesses strong data fitting and generalization capabilities, paving the way for new directions in PEFT development. However, there are still some issues that require further consideration and resolution:
1. While the Bone architecture shows better training performance, its peak memory usage is higher compared to LoRA and its variants.
2. It remains to be seen whether Bone is suitable for large models in other modalities.
We welcome the community to provide additional suggestions and conduct further tests.

### v2

@@
 Abstract
-As Large Language Models (LLMs) continue to grow in size, their computational and memory requirements increase correspondingly.
-Consequently, the exploration of cost-effective and efficient fine-tuning methods has become increasingly important.
 Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs.
 In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA.
-However, these methods also make the fine-tuning initialization process more complex, and it remains challenging to surpass the performance ceiling of full fine-tuning.
-To address these issues, this paper introduces an innovative method called Bone (Block Affine), which not only reduces memory overhead but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting.
+However, these improvements complicate the initial setup of model training and increase initialization time.
+More importantly, they overlook the internal interactions of the original weight information.
+To address these issues, we introduce a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information.
+Based on this theory, we designed a new PEFT technique called Bone (Block Affine), which not only enhances the utilization of original weight information but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting.
 Experimental comparisons across two different LLM architectures (LLaMA2, RWKV6) and various parameter scales demonstrate that the Bone structure can achieve rapid convergence and superior data fitting without the need for complex initialization.
@@
 However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios.
-Full-scale fine-tuning of large models is extremely costly; for example, training the LLaMA2-7B Touvron et al.
+Full-scale fine-tuning of large models is computationally costly; for example, finetuning the LLaMA2-7B Touvron et al.
 (2023) model with bfloat16 Wang & Kanwar (2019) precision requires around 60GB of VRAM.
-In contrast, by using PEFT (Parameter-Efficient Fine-Tuning)Xu et al.
-(2023) techniques, fine-tuning can be done with just a single 24GB VRAM GPU.
+In contrast, PEFT (Parameter-Efficient Fine-Tuning)Xu et al.
+(2023) techniques could reduce the VRAM requirement to fit into a 24GB VRAM GPU.
 As a result, numerous PEFT techniques and quantization methods have emerged to reduce the training costs of large models.
 LoRA (Low-Rank Adaptation) Hu et al.
-(2021) has become one of the most popular PEFT methods due to its small parameter size, its ability to optimize the model’s main weights, and the fact that it doesn’t add extra inference overhead.
+(2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after finetuning.
 LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$.
@@
 For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
-In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants and surpass full fine-tuning.
-It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights.
-To calculate the information interactions within the weight matrix $W$, Bone employs an initially zeroed, trainable matrix, named bone, to perform Block Affine operations on $W$.
+In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants.
+It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance.
+To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process.
+As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components.
+In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights.
+As the training steps increase, the influence of the original weights becomes more pronounced.
+This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
+To promote the internal interactions of information within the original weight matrix $W$, Bone employs an trainable zero-initiated matrix, named bone, to perform Block Affine operations on $W$.
 The specific calculations of Block Affine are detailed in Section 3, and the structural visualization is shown in Figure  2.
@@
 (2024)).
-As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  1, Table  2).
-Advantages of Bone::
-1. Bone does not require the complex initialization needed by LoRA variants, yet it achieves fast convergence and better data fitting.
-2. Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
-3. Bone provides a new direction for the development of PEFT techniques.
+As shown in Figure  3, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  2, Table  3).
+Our contributions can be summarized as follows
+1. We propose a new concept, “Weight Guide” which allows the continuous utilization of original weight information throughout the training process.
+2.
+We propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants.
+Block Affine not only efficiently promotes internal interactions of information within the weights but also enhances the weight utilization of low-rank matrices.
+3. The effectiveness of the Bone structure and “Weight Guide” has been validated through extensive experiments.
 (a) llama2-7b
-(b) RWKV6
+(b) Train Steps
 Figure 1: The left image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation.
-The right image displays the loss curves for RWKV6 with different parameter settings, also fine-tuned on MetaMathQA.
-Both models were trained with a batch size of 64 for a full epoch.
 Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings.
 Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
+The left image is a demonstration of the training state.
 2 Related Works
 The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al.
+(2019); He et al.
+(2022); Wang et al.
+(2022); Pfeiffer et al.
 (2020), prefix tuningLiu et al.
-(2023), prompt tuningBrown (2020), LoRAHu et al.
+(2023); Li & Liang (2021), prompt tuningBrown (2020); Liu et al.
+(2023); Lester et al.
+(2021); Razdaibiedina et al.
+(2023); Li & Liang (2021), LoRAHu et al.
+(2021); Meng et al.
+(2024); Wang et al.
+(2024); Si et al.
 (2024), and layer-freezing methods such as LISA.
@@
 Unlike prompts, prefixes are entirely composed of free parameters and do not correspond to actual tokens.
-Compared to traditional fine-tuning, prefix tuning only optimizes the prefix, meaning we only need to store a large Transformer model and a small set of task-specific prefixes, resulting in minimal overhead for additional tasks.
-Prompt tuning defines a unique prompt for each task, appending it to the input data while freezing the pre-trained model during training.
-LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers, while keeping the pre-trained weights unchanged.
+Compared to traditional fine-tuning, prefix tuning only optimizes a set of prefixes related to specific tasks, without any change to the original model.
+Prompt tuning defines a unique prompt for each task, prepending it to the input data while freezing the pre-trained model during training.
+LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers while keeping the pre-trained weights unchanged.
 This approach significantly addresses the inference latency introduced by adapters and the fitting challenges of prefix tuning.
 In LoRA, the adapter matrices $A$ and $B$ are updated with the same learning rate, but using the same rate for both may not effectively learn the features.
-LoRA+ corrects this suboptimality by setting different learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency.
+LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency.
 The DoRA Liu et al.
@@
 3 Bone: Block Affine
-In this section, we will introduce the Bone method.
-Unlike the LoRA series of architectures, Bone employs a single learnable matrix, initialized to zero and named ”bone.” This matrix interacts with the model’s matrix $W$ through Block Affine operations to represent $\Delta W$.
-This approach ensures that the initial state remains unchanged while also facilitating the integration of information between the weights.
-Assuming the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$.
-The specific calculation for Block Affine is as follows:
+In this section, we will introduce the Bone (Block Affine) method.
+Our design is based on multiple considerations.
+First, to achieve the aforementioned “Weight Guide” We modify the formula $Y=X(W+\Delta W)$ by replacing $\Delta W$ with $W\Delta W+\Delta W$, allowing the original weights to continuously guide (constrain) the direction of $\Delta W$ throughout the training process.
+To verify whether the guidance (constraint) from the original weights has a positive effect on the trainable matrix, we conducted an ablation experiment in Section  4.5, we named the comparison structure “Bone-Unconstrained”, with the comparison formula: $\Delta W=WA+B$.
+The experimental results indicate that “Weight Guide” not only provides positive guidance but also enhances the utilization of matrix parameters.
+Second, to enhance the fusion of internal weight features and extract more effective information, we opted for matrix multiplication over the Hadamard product in the $W\Delta W$ operation.
+The Hadamard product only captures features at corresponding positions, whereas matrix multiplication allows the accumulation of features from an entire row into a single feature.
+To validate this approach, we conducted experiments in Section  4.5 to compare different computational methods in terms of feature fusion and extraction.
+Here, we named the comparison structure “Bone-Hadamard” to facilitate readers’ understanding and reference in the following sections.
+Ultimately, we chose matrix multiplication as the core operation.
+Finally, we hypothesize that the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$. The specific calculation for Block Affine is as follows:
 $Y=X(W+\Delta W)$ (1)
 $\Delta W_{n,n}=W_{n,n}bone_{n,n}+bone_{n,n}$ (2)
-Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
-We conducted ablation experiments to verify the effectiveness of using a single learnable matrix structure.
-The comparison formulas are as follows:
-$\Delta W_{n,n}=W_{n,n}bone\_A_{n,n}+bone\_B_{n,n}$ (3)
-We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns are not equal.
-Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principles of fine-tuning.
+We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns is not equal.
+Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principle of parameter-efficiency.
 Therefore, we adjust $bone$ to be a low-rank matrix.
@@
 The final formula can be expressed as:
-$W_{gn,g,b,b}=Reshape(W_{n,m})$ (4)
+$W_{gn,g,b,b}=Reshape(W_{gn*b,g*b})=Reshape(W_{n,m})$ (3)
 $\Delta W_{n,m}=Reshape(W_{gn,g,b,b}bone_{g,b,b}+bone_{g,b,b})$ (4)
 During our comparative experiments between RWKV6 and LLaMA2, we found that the matrix $W$ has several different dimensions.
-For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$.
+For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp layer is $(4096,11008)$.
 These differences require us to continuously adjust g to ensure the structure functions correctly.
-To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, we modified the dimensions of Bone to $(n,b)$, where $b$ remains the size of the smallest block, and $n$ corresponds to the $n$ dimension of $W$.
-Thus, $g=n/b$.
+To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, We set $g=n/b$ in $bone_{g,b,b}$ allowing the parameters of Bone to be equivalent to those of $LoraA_{n,r}$.
 During computation, Bone only requires a dimension transformation, The Bone structure is illustrated in Figure 2(c), and the modified formula is as follows:
@@
 4 Experiments
-The experiments were conducted on 4×NVIDIA 4090 24G GPUs.In our experiments, we fixed lora_r=32 to ensure consistent trainable parameters across different fine-tuning methods, using the AdamW optimizer with a batch size of 64, a learning rate of 2e-5 and a warmup ratio of 0.0, without any weight decay.
+The experiments were conducted on 4×NVIDIA 4090 24G GPUs. Our experimental training parameter settings are shown in the table  1.
+Table 1: Training parameter settings
+Deivce Model low rank gobal batchsize learning rate precision
+4×NVIDIA 4090 RWKV6 r=32/b=64 64 2e-5 bf16
+Llama r=36/b=64 64 2e-5 bf16
 4.1 Evaluating the Performance of Bone under Different Parameter Settings in RWKV6
-Table  1presents a comparative evaluation of the fine-tuned RWKV6 models with different parameter settings on the NLG task.
-I fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al.
-(2023) and evaluated them on GSM8k Cobbe et al.
-(2021) and MATH datasets Yu et al.
-(2023).
-To ensure the accuracy of the experiment, the comparison between LoRA, PiSSA, and Bone was conducted with the same number of trainable parameters.
-The fine-tuning experiments for the 1.6B, 3B, and 7B RWKV6 models were each conducted for a complete epoch with a total batch size of 64 and a fixed learning rate of 2e-5.
+Models and Datasets
+To validate the stability of Bone, we trained RWKV6 models with different parameter sizes on math tasks.
+Math: We fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al. (2023) and evaluated them on GSM8k Cobbe et al. (2021) and MATH datasets Yu et al. (2023).
+Implementation Details
+For fine-tuning RWKV6 models with different parameter sizes, we kept the learning rate, batch size, and training steps the same across all experiments.
+We set lora_r=32 and bone_b=64 to ensure that the trainable parameters are equivalent between the two methods.
+Result
+As shown in Table 2, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning.
+For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes.
+This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
+(a) RWKV6 loss
+Figure 3: The image displays the loss curves for RWKV6 with different parameter settings, fine-tuned on MetaMathQA.
 Table 2: The performance of different PEFT techniques under various parameter settings on RWKV6.
@@
 Bone 88.1M 42.76 6.34
-As shown in Table 1, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning.
-For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes.
-This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
 4.2 Experiments on Llama2-7B
 Models and Datasets
-[…]
+To verify the generalizability of Bone, we conducted more comprehensive experiments on LLaMA2-7B. we conducted 3 more task finetuning experiments on LLaMa2-7B: math, code, and chat.
 1.
-Math: We train our model on a 395k subset of MetaMathQA Yu et al.
+Math: We trained our model on a 395k subset of MetaMathQA Yu et al.
 (2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al.
@@
 (2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects.
-The quality of the responses is judged by GPT-4o, and we report the first turn score.
+We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
+Implementation Details
+[…]
 Result
-The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other structures in terms of convergence speed, data fitting, and generalization capabilities.
-Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further demonstrating the generalizability of the Bone approach.
+The results, as shown in Table  3 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
+Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further Bone outperformed methods compared, on all 3 different tasks focusing on various disciplines, demonstrating the generalizability.
 Table 3: We fine-tuned LLaMA2-7B using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
@@
 Bone 87.0M 49.36 8.88 25.62 4.81
-4.3 Experiments were conducted with different values of block_size $(b)$.
+4.3 Experiments on different values of block_size $(b)$.
 This subsection explores the upper limits of the Bone structure by varying the block_size in the Bone matrix.
@@
 Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks.
-The test results, as shown in Table  3, demonstrate that the fine-tuning performance improves as the value of b increases.
+The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases.
 Notably, when b reaches 16, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark.
-However, its performance on the Math benchmark is weaker, indicating that the value of b plays a crucial role in determining the model’s generalization ability.
-As b increases, the model’s generalization capacity strengthens.
+However, its performance on the Math benchmark is only 3.54.
+The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of b impacts the model’s ability to understand unseen data.
+Based on this observation, we conclude that increasing b not only raises the number of parameters but also enlarges the window for information fusion, allowing the model to perform better on unseen data.
 Table 4: Comparing different values of block_size $(b)$
@@
 We fine-tuned LLaMA2-7B on the MetaMathQA dataset and validated the results on GSM8k and Math.
-The results are shown in Table  4.
+The results are shown in Table  5.
 Since in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bone structure, making it difficult to align parameter counts.
@@
 This allows for a fairer comparison between Bone-row and Bone-col.
-The experimental results, shown in Table  4, indicate that the difference between the two is minimal, with both performing well.
+The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well.
 Therefore, we believe that Bone can effectively fit data regardless of the grouping method used.
@@
 Bone-col 87.0M 49.36 8.88
-RWKV6-3B Bone-row 44.1M 25.93 3.32
-Bone-col 44.1M 25.25 2.82
+RWKV6-3B Bone-row 55.1M 25.93 3.32
+Bone-col 55.1M 25.25 2.82
 4.5 Ablation experiments on Block Affine.
-In this subsection, we aim to verify the effectiveness of a single learnable matrix, and we name the comparison structure Bone-AB.
-It consists of two low-rank matrices and is grouped by rows.
-To ensure the trainable parameter count is consistent, we set $b=32$ in Bone-AB and $b=64$ in Bone-row.
-For this comparison experiment, we continue to evaluate the Math task, with the results shown in Table  5.
-Bone-AB performed the worst, further demonstrating Block Affine not only efficiently promotes information exchange between weights but also enhances the weight utilization of low-rank matrices.
+We will conduct ablation experiments on the Block Affine structure, which consist of two parts:
+1.
+Verifying the effectiveness of “Weight Guide” on LLaMA2-7B: The results are shown in the table  6.
+It is evident that the performance of Bone-Unconstrained is even worse than that of Bone-Col $b=32$, which has half the number of parameters but incorporates the “Weight Guide” characteristic.
+2. Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
+The comparison formulas are as follows:
+$Bone-Unconstrained\Delta W_{n,n}=W_{n,n}A_{n,n}+B_{n,n}$ (11)
+$Bone-Hadamard\Delta W_{n,n}=W_{n,n}*bone_{n,n}+bone_{n,n}$ (12)
 Table 6: Block Affine on math tasks
@@
 Bone-row b=64 72.8M 45.76 7.82
-Bone-AB b=32 72.8M 44.35 6.16
+Bone-Unconstrained b=32 72.8M 44.35 6.16
+RWKV6-3B Bone 55.1M 25.93 3.32
+Bone-Hadamard 55.1M 22.44 2.67
 5 Resource and efficiency
-Table  6 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU.
+Table  7 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU.
 The specific fine-tuning settings are as follows: batchsize = 1, context length (ctx_len) = 512.
-Due to the structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage by 5GB compared to LoRA.
-Additionally, Bone’s token throughput is only two-thirds of LoRA’s.
+The structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage.
 To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput.
-Despite the fact that the Bone structure significantly outperforms LoRA variants in fine-tuning, we cannot ignore the efficiency issues.
+At the end of the table, we provide the real-world training scenarios comparing the memory usage and token throughput of Bone and LoRA.
 Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
@@
 Bone(checkpoint) 33.0M 7398 MB 3.46 kt/s
+Bone(model checkpoint) 33.0M 7398 MB 1.59 kt/s
 RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
@@
 Bone(checkpoint) 55.1M 12168 MB 1.72 kt/s
+RWKV-3B LoRA(model checkpoint) 55.1M 4*15328 MB 15.6 kt/s
+Bone(model checkpoint) 55.1M 4*15305 MB 14.2 kt/s
 6 Conclusion
-This paper introduces a new PEFT technique called Bone(Block Affine), which employs block affine transformations on the weight matrix to capture the internal correlations within $W$, while also avoiding the complexity of initialization.
+This paper introduces a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information.
+Based on this theory, we designed a new PEFT technique called Bone (Block Affine).
+Additionally, we employed block matrix multiplication to efficiently promote information fusion within the weights in the Bone structure.
 Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA.

### v4

@@
-Bone: Block Affine Transformation as Parameter Efficient Fine-tuning Methods for Large Language Models
+Bone: Block Affine Adaptation of Large Language Models
 Abstract
@@
 In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA.
-However, these improvements complicate the initial setup of model training and increase initialization time.
-More importantly, they overlook the internal interactions of the original weight information.
-To address these issues, we introduce a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information.
-Based on this theory, we designed a new PEFT technique called Bone (Block Affine), which not only enhances the utilization of original weight information but also emphasizes the internal connections between weights, leading to faster convergence and better data fitting.
-Experimental comparisons across two different LLM architectures (LLaMA2, RWKV6) and various parameter scales demonstrate that the Bone structure can achieve rapid convergence and superior data fitting without the need for complex initialization.
-For example, when fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8k and math benchmarks, Bone achieved fine-tuning scores of 49.36 and 8.8, respectively, outperforming PISSA by 5.84% and 1.96%.
+This paper introduces a novel PEFT technique distinct from LoRA, called Block-Affine Adaptation (Bone).
+By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants.
+Compared to LoRA, Bone significantly reduces memory usage and achieves faster computation.
+Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants.
+Inspired by Pissa, we further proposed the “Weight Guide” theory to better utilize the information from the original weights.
+By integrating “Weight Guide” with Bone, we developed a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
 1 Introduction
@@
 LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$.
-Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The Figure 2(a) illustrates the structure visualization.
+Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The figure  2 illustrates the structure visualization.
 However, extensive experiments Ding et al.
@@
 For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
-In this paper, we propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants.
-It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance.
-To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process.
-As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components.
-In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights.
-As the training steps increase, the influence of the original weights becomes more pronounced.
-This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
-To promote the internal interactions of information within the original weight matrix $W$, Bone employs an trainable zero-initiated matrix, named bone, to perform Block Affine operations on $W$.
-The specific calculations of Block Affine are detailed in Section 3, and the structural visualization is shown in Figure  2.
+As research on LoRA becomes increasingly saturated, it is essential to explore new directions for PEFT techniques and effectively leverage the unique characteristics of LLMs.
+In this work, we first designed Block-Affine Adaptation (Bone), which divides the original weights into multiple subspaces that share a single trainable matrix for updates, making it more efficient than LoRA’s two low-rank matrices.
+Secondly, to address the limitations of LoRA and fully utilize the information in the original weights, we propose a new theory, “Weight Guide”.
+“Weight Guide” enables feature extraction by applying a simple linear transformation to the original weights.
+Finally, to validate the effectiveness of “Weight Guide”, we combined Bone with “Weight Guide” to develop Block-Affine Transformation (Bat).
+Our extensive evaluation shows that Bone not only excels in both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks but also significantly outperforms LoRA and its variants.
+Additionally, Bone retains the advantages of LoRA, such as ease of use and no additional computational overhead for the model.
+In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board.
 To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al.
@@
 (2024)).
-As shown in Figure  3, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities (Table  2, Table  3).
+As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities ( Table  2).
+As a completely new structure distinct from LoRA, Bone offers clear improvements in computational efficiency and memory savings, as detailed in table  7.
 Our contributions can be summarized as follows
-1. We propose a new concept, “Weight Guide” which allows the continuous utilization of original weight information throughout the training process.
+1.
+We propose a novel PEFT technique called Block-Affine Adaptation (Bone).
+Bone outperforms LoRA and its variants across various fine-tuning tasks.
+Additionally, Bone is more memory-efficient and computationally faster compared to LoRA.
 2.
-We propose a novel PEFT method called Bone (Block Affine) to overcome the limitations of LoRA variants.
-Block Affine not only efficiently promotes internal interactions of information within the weights but also enhances the weight utilization of low-rank matrices.
-3. The effectiveness of the Bone structure and “Weight Guide” has been validated through extensive experiments.
+To effectively leverage the implicit information in the original weights, we propose the “Weight Guide” theory.
+By integrating this theory with Bone, we design Block-Affine Transformation (Bat), demonstrating the effectiveness of “Weight Guide”.
 (a) llama2-7b
@@
 LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
-3 Bone: Block Affine
-In this section, we will introduce the Bone (Block Affine) method.
-Our design is based on multiple considerations.
-First, to achieve the aforementioned “Weight Guide” We modify the formula $Y=X(W+\Delta W)$ by replacing $\Delta W$ with $W\Delta W+\Delta W$, allowing the original weights to continuously guide (constrain) the direction of $\Delta W$ throughout the training process.
-To verify whether the guidance (constraint) from the original weights has a positive effect on the trainable matrix, we conducted an ablation experiment in Section  4.5, we named the comparison structure “Bone-Unconstrained”, with the comparison formula: $\Delta W=WA+B$.
-The experimental results indicate that “Weight Guide” not only provides positive guidance but also enhances the utilization of matrix parameters.
-Second, to enhance the fusion of internal weight features and extract more effective information, we opted for matrix multiplication over the Hadamard product in the $W\Delta W$ operation.
-The Hadamard product only captures features at corresponding positions, whereas matrix multiplication allows the accumulation of features from an entire row into a single feature.
-To validate this approach, we conducted experiments in Section  4.5 to compare different computational methods in terms of feature fusion and extraction.
-Here, we named the comparison structure “Bone-Hadamard” to facilitate readers’ understanding and reference in the following sections.
-Ultimately, we chose matrix multiplication as the core operation.
-Finally, we hypothesize that the matrix $W$ has dimensions $(n,n)$, ideally, we train with a full-rank matrix $bone$ $(n,n)$. The specific calculation for Block Affine is as follows:
-$Y=X(W+\Delta W)$ (1)
-$\Delta W_{n,n}=W_{n,n}bone_{n,n}+bone_{n,n}$ (2)
-We know that in LLMs, the matrix $W$ is typically very large, and in most cases, the number of rows and columns is not equal.
-Setting $bone$ to the same dimensions as W would result in excessive computational overhead, which goes against the principle of parameter-efficiency.
-Therefore, we adjust $bone$ to be a low-rank matrix.
-When the weight matrix $W$ has dimensions $(n,m)$, it is divided according to a block_size $b$ into $W(n/b,b,m/b,b)$, yielding $(n/b)\times(m/b)$ blocks.
-Inspired by GQA(Grouped Query Attention)Ainslie et al.
-(2023), we group these blocks together to significantly reduce the number of trainable parameters without affecting performance.
-Different groups share a single Bone matrix for computation.
-Thus, the Bone matrix is set to $(g,b,b)$, where $g$ is the group size.
-The $W$ matrix is reshaped to $(gn,g,b,b)$, where $gn$ represents the number of groups.
-The Bone structure is illustrated in Figure 2(b).
-The final formula can be expressed as:
-$W_{gn,g,b,b}=Reshape(W_{gn*b,g*b})=Reshape(W_{n,m})$ (3)
-$\Delta W_{n,m}=Reshape(W_{gn,g,b,b}bone_{g,b,b}+bone_{g,b,b})$ (4)
-During our comparative experiments between RWKV6 and LLaMA2, we found that the matrix $W$ has several different dimensions.
-For example, in RWKV6-7B, the dimension of ffn’s $k$ is $(4096*3.5,4096)$, while in LLaMA2-7B, the dimension of gate_proj in the mlp layer is $(4096,11008)$.
-These differences require us to continuously adjust g to ensure the structure functions correctly.
-To better compare Bone with various LoRA variants and to address the compatibility issues arising from differences in FFN layers across different models and parameter settings, We set $g=n/b$ in $bone_{g,b,b}$ allowing the parameters of Bone to be equivalent to those of $LoraA_{n,r}$.
-During computation, Bone only requires a dimension transformation, The Bone structure is illustrated in Figure 2(c), and the modified formula is as follows:
-$W_{m/b,n/b,b,b}=Reshape(W_{n,m})$ (5)
-$bone_{n/b,b,b}=Reshape(bone_{n,b})$ (6)
-$\Delta W_{n,m}=Reshape(W_{m/b,n/b,b,b}bone_{n/b,b,b}+bone_{n/b,b,b})$ (7)
-In this way, we can easily adapt to any dimension of W and models with different architectures.
-The compatibility method used above groups by columns, and naturally, we can also group by rows.
-The visualization is shown in Figure  2(d), and the formula is as follows:
-$W_{n/b,m/b,b,b}=Reshape(W_{n,m})$ (8)
-$bone_{m/b,b,b}=Reshape(bone_{m,b})$ (9)
-$\Delta W_{n,m}=Reshape(W_{n/b,m/b,b,b}bone_{m/b,b,b}+bone_{m/b,b,b})$ (10)
-(a) LoRA
-(b) Bone-both
-(c) Bone-col
-(d) Bone-row
+3 Method
+3.1 Motivation
+The primary motivations are as follows:
+Firstly, current research on LoRA has reached a saturation point, underscoring the need for a new direction in Parameter-Efficient Fine-Tuning (PEFT) techniques.
+Secondly, existing PEFT methods lack a thorough exploration and effective utilization of the original model weights.
+3.2 Bone: Block-Affine update matrices
+To design a novel PEFT structure that rivals LoRA, we conducted an in-depth study of dense matrices.
+Current theories suggest that LLM weights exhibit sparse properties.
+For example, LoRA utilizes low-rank matrices to update dense matrices, while techniques like GQA Ainslie et al.
+(2023) and MQA Shazeer (2019) employ shared mechanisms to reduce parameter overhead with remarkable effectiveness.
+Inspired by these approaches, we hypothesize that dividing LLM weights into multiple subspaces allows different subspaces to share a single low-rank matrix for updates.
+As illustrated in the figure  2, we name this structure Block-Affine Adaptation (Bone).
+It is evident that the ways to partition LLM weights are diverse, as illustrated in the figure.
+Different partitioning methods influence the shape of the Bone matrix, the computation logic, and ultimately, the processing speed.
+In this work, we propose the Best Bone structure.
+For a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$, we update it using a single matrix: $W+\Delta W=W+expand(B)$, where $B\in\mathbb{R}^{r\times d}$, and the rank $r\ll k$.
+This design not only retains the simplicity and ease of use characteristic of LoRA but also achieves higher computational efficiency.
+The calculation is expressed as follows:
+$Y=Wx+\Delta Wx=Wx+sum(x)B$ (1)
+(a) Bone instruciton
+Figure 2: Bone requires training only a single low-rank matrix initialized to zero. Best Bone represents the optimal design structure of Bone. Subsequent experiments default to using Best Bone.
+3.3 “Weight Guide”
+It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance.
+To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process.
+As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components.
+In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights.
+As the training steps increase, the influence of the original weights becomes more pronounced.
+This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
+So, how can we extract the critical information from the original weights?
+A straightforward approach is to apply a learnable matrix directly to the original weights using the Hadamard product.
+This method essentially assigns a weight to each element, allowing the learnable matrix to autonomously determine the importance of individual elements.
+Although this method is simple and straightforward, the elements within the weights remain isolated, making it impossible to learn interactions between them.
+Therefore, we replace the Hadamard product with matrix multiplication, essentially applying a simple linear transformation.
+This approach links the entire weight matrix space, providing a better chance to discover the optimal solution.
+The corresponding formula is shown below:
+$\Delta W_{n,n}$ $=W_{n,n}\odot bone_{n,n}$ (2)
+$\Delta W_{n,n}$ $=W_{n,n}\otimes bone_{n,n}$
+3.4 Bat: Block-Affine Transformation
+In the previous section, we introduced Bone, which is both simple and efficient.
+However, applying the same updates to different subspaces of the weight matrix $W\in\mathbb{R}^{d\times k}$ is clearly counterintuitive.
+To allow for more flexibility in the update directions of different subspaces, we combine Bone with “Weight Guide”, proposing a new method called Block-Affine Transformation (Bat).
+To balance computational efficiency, we divide the original weights into smaller subspaces compared to Bone.
+Here, we name the trainable matrix $bone\in\mathbb{R}^{r\times d}$, where its rank $r$ is replaced with block size $b$.
+The formula for Bone-col is as follows:
+$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (3)
+$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
+$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\otimes bone_{d/b,b,b}+bone_{d/b,b,b})$
+Similar to Bone, Bat can also adopt various grouping methods, as illustrated in the figure 3.
+In Section 6, we compared the performance of different partitioning strategies for the structure.
+The formula for Bone-row is as follows:
+$W_{d/b,k/b,b,b}$ $=Reshape(W_{d,k})$ (4)
+$bone_{k/b,b,b}$ $=Reshape(bone_{b,k})$
+$\Delta W_{d,k}$ $=Reshape(W_{d/b,k/b,b,b}\otimes bone_{k/b,b,b}+bone_{k/b,b,b})$
+(a) Bone-both
+(b) Bone-col
+(c) Bone-row
 […]
 4 Experiments
-The experiments were conducted on 4×NVIDIA 4090 24G GPUs. Our experimental training parameter settings are shown in the table  1.
-Table 1: Training parameter settings
-Deivce Model low rank gobal batchsize learning rate precision
-4×NVIDIA 4090 RWKV6 r=32/b=64 64 2e-5 bf16
-Llama r=36/b=64 64 2e-5 bf16
-4.1 Evaluating the Performance of Bone under Different Parameter Settings in RWKV6
+In this section, we evaluate the performance of Bone on various benchmark datasets.
+Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model.
+Subsequently, we evaluated the Natural Language Generation (NLG) capabilities by fine-tuning the LLM.
+The experiments were conducted on 4×NVIDIA 4090 24G GPUs.
+4.1 Experiments on Natural Language Understanding
 Models and Datasets
-To validate the stability of Bone, we trained RWKV6 models with different parameter sizes on math tasks.
-Math: We fine-tuned RWKV6 models with varying parameters on the MetaMathQA dataset Yu et al. (2023) and evaluated them on GSM8k Cobbe et al. (2021) and MATH datasets Yu et al. (2023).
+We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC.
+Performance is evaluated on the development set using accuracy as the primary metric.
 Implementation Details
-For fine-tuning RWKV6 models with different parameter sizes, we kept the learning rate, batch size, and training steps the same across all experiments.
-We set lora_r=32 and bone_b=64 to ensure that the trainable parameters are equivalent between the two methods.
-Result
-As shown in Table 2, Bone consistently outperforms PiSSA and LoRA across all parameter settings and tasks during fine-tuning.
-For example, on the GSM8k test, Bone achieved improvements of 1.82%, 0.61%, and 2.28% over PiSSA across different model sizes.
-This further demonstrates that the Bone architecture is adaptable to models with various parameter settings and consistently surpasses other fine-tuning structures.
-(a) RWKV6 loss
-Figure 3: The image displays the loss curves for RWKV6 with different parameter settings, fine-tuned on MetaMathQA.
-Table 2: The performance of different PEFT techniques under various parameter settings on RWKV6.
-Model Strategy Trainable GSM8K Math
-RWKV 6-1.6B LoRA 33.0M 10.84 1.14
-PiSSA 33.0M 12.89 1.28
-Bone 33.0M 14.71 1.94
-RWKV 6-3B LoRA 55.1M 21.22 2.34
-PiSSA 55.1M 24.64 2.54
-Bone 55.1M 25.25 2.82
-RWKV 6-7B LoRA 88.1M 38.13 6.06
-PiSSA 88.1M 40.48 6.12
-Bone 88.1M 42.76 6.34
-4.2 Experiments on Llama2-7B
+The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU.
+Each experiment is conducted with 3 different random seeds, and the average performance is reported.
+Results
+As shown in Table 1, Bone demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and Pissa.
+Table 1: The results of fine-tuning RoBERTa-base using Bone and various LoRA variants were compared on a subset of the GLUE benchmark.
+Method Trainable MNLI SST-2 CoLA QNLI MRPC
+LoRA 0.2365% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
+Pissa 0.2365% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
+Bone 0.2360% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
+4.2 Experiment on Natural Language Generation
 Models and Datasets
-To verify the generalizability of Bone, we conducted more comprehensive experiments on LLaMA2-7B. we conducted 3 more task finetuning experiments on LLaMa2-7B: math, code, and chat.
+To verify the generalizability of Bone, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math, code, and chat.
 1.
@@
 Implementation Details
-[…]
+The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
+It is worth noting that the weights of LLaMA2-7B are not fully symmetric, making it impossible to perfectly align the trainable parameters when comparing Bone and LoRA.
+To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Bone to 64, ensuring that Bone uses fewer parameters than LoRA to demonstrate its superiority.
+Each experiment is conducted with 2 different random seeds, and the average performance is reported.
 Result
-The results, as shown in Table  3 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
-Specifically, Bone outperforms PISSA by 5.84% on GSM8K and by 4.89% on HumanEval, further Bone outperformed methods compared, on all 3 different tasks focusing on various disciplines, demonstrating the generalizability.
-Table 3: We fine-tuned LLaMA2-7B using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
+The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
+Bone demonstrates outstanding performance across three different tasks.
+On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants.
+On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
+Table 2: We fine-tuned LLMs using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
 Model Strategy Trainable GSM8K Math HumanEval MT-Bench
-Llama2-7B LoRA 89.9M 38.97 0.44 16.46 3.93
-OLoRA 89.9M 41.93 6.48 20.12 4.09
-PiSSA 89.9M 43.52 6.92 20.73 4.25
-Bone 87.0M 49.36 8.88 25.62 4.81
-4.3 Experiments on different values of block_size $(b)$.
-This subsection explores the upper limits of the Bone structure by varying the block_size in the Bone matrix.
-Notably, while we change the size of $b$, the value of $n$ remains constant.
-Therefore, as $b$ increases, the dimensions of Bone transform into the bone_g form, leading to a corresponding reduction in the number of blocks within each group.
+Llama2-7B LoRA 89.9M 40.75 5.22 17.68 3.73
+OLoRA 89.9M 42.93 6.51 21.12 4.03
+PiSSA 89.9M 43.89 6.92 22.25 4.11
+Bone 87.0M 48.16 8.58 24.08 4.31
+RWKV 6-7B LoRA 88.1M 38.13 6.06 - -
+PiSSA 88.1M 40.48 6.12 - -
+Bone 88.1M 41.73 6.52 - -
+4.3 Effect of Rank $r$
+This subsection explores the upper limits of the Bone structure by varying the rank $r$ in the Bone matrix.
 Comparative experiments were conducted by fine-tuning LLaMA2-7B on the MetaMathQA dataset and validating on GSM8K and Math benchmarks.
-The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases.
-Notably, when b reaches 16, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark.
-However, its performance on the Math benchmark is only 3.54.
-The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of b impacts the model’s ability to understand unseen data.
-Based on this observation, we conclude that increasing b not only raises the number of parameters but also enlarges the window for information fusion, allowing the model to perform better on unseen data.
-Table 4: Comparing different values of block_size $(b)$
+The test results, as shown in Table  3, demonstrate that the fine-tuning performance improves as the value of b increases.
+Notably, when $r=16$, the Bone structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark.
+However, its performance on the Math benchmark is only 3.73.
+The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data.
+Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
+Table 3: Comparing different values of rank $(r)$
+Model rank Trainable GSM8K Math
+Llama2-7B 16 21.7M 45.90 3.77
+32 43.5M 46.18 7.43
+64 87.0M 48.16 8.58
+128 174.0M 53.49 10.08
+Figure 4: Training loss curves of Bone with different rank $r$ on the MetaMathQA dataset.
+4.4 Bone vs Bat
+To validate the effectiveness of the “Weight Guide”, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K.
+As shown in the table, Bat, equipped with “Weight Guide”, achieves significant improvements in performance metrics compared to Bone.
+Table 4: Comparing Bat-row, Bat-col on math tasks
 Model Strategy Trainable GSM8K Math
-Llama2-7B LoRA 89.9M 38.97 0.44
-PiSSA 89.9M 43.52 6.92
-Bone b=16 21.7M 45.79 3.54
-Bone b=32 43.5M 46.70 7.67
-Bone b=64 87.0M 49.36 8.88
-Bone b=128 174.0M 53.15 10.56
-Figure 4: Training loss curves of Bone with different block_size on the MetaMathQA dataset.
-4.4 Ablation experiments on different grouping methods.
-In this subsection, we explore the impact of different grouping methods in the Bone structure on model fine-tuning performance.
-Due to structural differences in the weight matrix W, the Bone-free grouping requires manual configuration, which is inconvenient.
+Llama2-7B Bone 87.0M 48.16 8.58
+Bat 87.0M 49.36 8.88
+RWKV6-3B Bone 55.1M 41.73 6.52
+Bat 55.1M 42.76 6.60
+4.5 Ablation experiments on different grouping methods for Bat
+In this subsection, we explore the impact of different grouping methods in the Bat structure on model fine-tuning performance.
+Due to structural differences in the weight matrix W, the Bat-free grouping requires manual configuration, which is inconvenient.
 Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure.
@@
 The results are shown in Table  5.
-Since in LLaMA2-7B, the dimension of gate_proj in the mlp part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bone structure, making it difficult to align parameter counts.
-Even though Bone-row performs better than Pissa with only around 20M parameters, the gap between Bone-row and Bone-col is too large.
-Initially, we suspected this gap was due to Bone-row lacking about 14M parameters compared to Bone-col.
-However, this assumption was disproved when we observed that Bone-row with 32 blocks and 43.5M parameters performed similarly to Bone-row with 72.8M parameters.
-This made it difficult to directly evaluate which grouping method performs better.
-To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bone fine-tuning.
-This allows for a fairer comparison between Bone-row and Bone-col.
+Since in LLaMA2-7B, the dimension of gate_proj in the MLP part is $(4096,11008)$, this leads to an asymmetry between row-wise and column-wise grouping in the Bat structure, making it difficult to align parameter counts.
+Although Bat-row uses 15M fewer parameters than Bat-col, it still delivers excellent performance.
+However, this discrepancy in parameter counts makes it challenging to accurately evaluate the differences between the grouping methods.
+To explore the differences between the two grouping methods and the effect of parameter count, we added a comparative experiment with RWKV6-3B, as RWKV6’s symmetrical structure ensures that the trainable parameter count is the same whether using row-wise or column-wise grouping during Bat fine-tuning.
+This allows for a fairer comparison between Bat-row and Bat-col.
 The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well.
-Therefore, we believe that Bone can effectively fit data regardless of the grouping method used.
-The key factor influencing Bone’s performance remains the block size.
+Therefore, we believe that Bat can effectively fit data regardless of the grouping method used.
+The key factor influencing Bat’s performance remains the block size.
 However, this doesn’t imply that different grouping methods are meaningless.
@@
 (2024); Lee et al.
-(2024); Shazeer (2019), the main weight matrices become smaller, and Bone will need to adjust its grouping or employ other techniques to adapt to these new technologies.
-Table 5: Comparing Bone-row, Bone-col on math tasks
+(2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
+Table 5: Comparing Bat-row, Bat-col on math tasks
 Model Strategy Trainable GSM8K Math
-Llama2-7B Bone-row 72.8M 45.76 7.82
-Bone-col 87.0M 49.36 8.88
-RWKV6-3B Bone-row 55.1M 25.93 3.32
-Bone-col 55.1M 25.25 2.82
-4.5 Ablation experiments on Block Affine.
-We will conduct ablation experiments on the Block Affine structure, which consist of two parts:
-1.
-Verifying the effectiveness of “Weight Guide” on LLaMA2-7B: The results are shown in the table  6.
-It is evident that the performance of Bone-Unconstrained is even worse than that of Bone-Col $b=32$, which has half the number of parameters but incorporates the “Weight Guide” characteristic.
-2. Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
-The comparison formulas are as follows:
-$Bone-Unconstrained\Delta W_{n,n}=W_{n,n}A_{n,n}+B_{n,n}$ (11)
-$Bone-Hadamard\Delta W_{n,n}=W_{n,n}*bone_{n,n}+bone_{n,n}$ (12)
-Table 6: Block Affine on math tasks
+Llama2-7B Bat-row 72.8M 45.76 7.82
+Bat-col 87.0M 49.36 8.88
+RWKV6-3B Bat-row 55.1M 25.93 3.12
+Bat-col 55.1M 25.25 3.09
+4.6 Ablation experiments on the specific implementation of “Weight Guide”
+Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
+The formula for Bone-Hadamard is as follows:
+$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (5)
+$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
+$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\odot bone_{d/b,b,b}+bone_{d/b,b,b})$
+Table 6: Block-Affine on math tasks
 Model Strategy Trainable GSM8K Math
-Llama2-7B Bone-col b=32 43.5M 46.70 7.67
-Bone-row b=64 72.8M 45.76 7.82
-Bone-Unconstrained b=32 72.8M 44.35 6.16
-RWKV6-3B Bone 55.1M 25.93 3.32
-Bone-Hadamard 55.1M 22.44 2.67
+RWKV6-3B Bat 55.1M 25.93 3.32
+Bat-Hadamard 55.1M 22.44 2.67
 5 Resource and efficiency
-Table  7 compares the training resources and token throughput required for fine-tuning RWKV6-3B using Bone and LoRA on a single 4090 GPU.
+Table  7 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU.
 The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
-The structure of Bone, an additional intermediate variable for the W weight is created during training, which increases memory usage.
-To address this, we applied checkpointing specifically to the Bone structure, eliminating the extra intermediate variable, which successfully reduced memory usage but also further decreased token throughput.
-At the end of the table, we provide the real-world training scenarios comparing the memory usage and token throughput of Bone and LoRA.
-Therefore, a key focus of our future work will be improving the Bone operator to enhance token throughput and reduce memory usage.
+The results show that Bone has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient.
+However, Bat incurs significantly higher memory usage due to large intermediate values and is slower in comparison.
+At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
+Therefore, a key focus of our future work will be improving the Bat operator to enhance token throughput and reduce memory usage.
 Table 7: Resource and efficiency
 Model Strategy Trainable GPU Memory Token throughput
-RWKV-1.6B LoRA 33.0M 7314 MB 5.86 kt/s
-Bone 33.0M 12276 MB 4.11 kt/s
-Bone(checkpoint) 33.0M 7398 MB 3.46 kt/s
-Bone(model checkpoint) 33.0M 7398 MB 1.59 kt/s
 RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
-Bone 55.1M 22978 MB 2.16 kt/s
-Bone(checkpoint) 55.1M 12168 MB 1.72 kt/s
-RWKV-3B LoRA(model checkpoint) 55.1M 4*15328 MB 15.6 kt/s
-Bone(model checkpoint) 55.1M 4*15305 MB 14.2 kt/s
+Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.95}$ kt/s
+Bat 55.1M 22978 MB 2.16 kt/s
+RWKV-3B LoRA(use checkpoint) 55.1M 4*15328 MB 15.6 kt/s
+Bone(use checkpoint) 55.1M $\mathbf{4*15304}$ MB $\mathbf{16.0}$ kt/s
+Bat(use checkpoint) 55.1M 4*15305 MB 14.2 kt/s
 6 Conclusion
-This paper introduces a novel theory, “Weight Guide” aimed at continuously guiding trainable matrices through the original weights during training to enhance the utilization of weight information.
-Based on this theory, we designed a new PEFT technique called Bone (Block Affine).
-Additionally, we employed block matrix multiplication to efficiently promote information fusion within the weights in the Bone structure.
-Through extensive experiments, we found that Bone can achieve a rapid drop in loss during the early stages of training, similar to or even better than PISSA.
-Additionally, Bone demonstrates superior data fitting in the later stages of training and shows exceptional performance across different LLM architectures and parameter scales, proving its versatility.
+This work, inspired by GQA and MQA, leverages the sparsity of LLM weights to design the Block-Affine Adaptation (Bone) structure.
+In Bone, the original weights are divided into multiple subspaces, all of which share a single low-rank matrix initialized to zero for updates.
+Extensive experiments demonstrate that Bone consistently outperforms LoRA and its variants across various tasks, while also offering superior computational efficiency.
+To break the limitations of LoRA and effectively utilize the information from the original weights, we propose the “Weight Guide” theory, which enables significant improvements through simple linear transformations.
+By integrating these innovations, we introduce a new structure called Block-Affine Transformation (Bat).
+Bat not only validates the effectiveness of “Weight Guide” but also addresses the limitation of Bone, where identical updates are applied to all subspaces.
+Experimental results show that Bat surpasses Bone on multiple tasks, although with a trade-off in computational efficiency.
 Bone brings new possibilities to existing LLM PEFT techniques.
 Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
-7 Limitations and future work
-Due to limited training resources, this paper lacks comparative experiments on full fine-tuning.
-However, the extensive experiments presented demonstrate that Bone possesses strong data fitting and generalization capabilities, paving the way for new directions in PEFT development.
-However, there are still some issues that require further consideration and resolution:
-1. While the Bone architecture shows better training performance, its peak memory usage is higher compared to LoRA and its variants.
-2. It remains to be seen whether Bone is suitable for large models in other modalities.
+7 Future work
+Bone and “Weight Guide” are merely a starting point, offering a foundation for researchers to explore additional branches inspired by Bat. There are many directions for future work.
+1. Can Bone adapt to multimodal tasks and other complex scenarios?
+2. Can “Weight Guide” be implemented using alternative computation methods?
 We welcome the community to provide additional suggestions and conduct further tests.

### v5

@@
-Bone: Block Affine Adaptation of Large Language Models
+Bone: Block-Affine Adaptation of Large Language Models
 Abstract
 Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs.
-In pursuit of performance closer to full-parameter training, a series of LoRA variants have emerged, such as LoRA+, PISSA, Olora, and LoRA-GA.
-This paper introduces a novel PEFT technique distinct from LoRA, called Block-Affine Adaptation (Bone).
+Many LoRA variants have emerged, yet they lack a design tailored to the characteristics of LLM weights and fail to leverage the original weights effectively.
+To address the sparsity of LLM weights, and drawing inspiration from GQA and MQA, we propose Block-Affine Adaptation (Bone), a novel PEFT technique distinct from LoRA.
 By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants.
@@
 Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants.
-Inspired by Pissa, we further proposed the “Weight Guide” theory to better utilize the information from the original weights.
-By integrating “Weight Guide” with Bone, we developed a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
+Inspired by Pissa, we propose a new theory called “Weight Guide” to better utilize the information embedded in the original weights.
+This approach extracts valuable information through a linear transformation of the original weight matrix using a trainable matrix.
+To validate the effectiveness of “Weight Guide” we combined it with Bone to create a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
 1 Introduction
@@
 Method Trainable MNLI SST-2 CoLA QNLI MRPC
-LoRA 0.2365% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
-Pissa 0.2365% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
-Bone 0.2360% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
+LoRA 0.236% $85.63_{\pm 0.01}$ $\mathbf{94.03_{\pm 0.02}}$ $62.40_{\pm 0.71}$ $91.37_{\pm 0.97}$ $87.98_{\pm 0.23}$
+Pissa 0.236% $\mathbf{85.72_{\pm 0.40}}$ $93.64_{\pm 0.13}$ $67.28_{\pm 0.59}$ $91.40_{\pm 0.54}$ $88.11_{\pm 0.24}$
+Bone 0.236% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
 4.2 Experiment on Natural Language Generation
@@
 Bone 88.1M 41.73 6.52 - -
+Mistral-7B LoRA 89.1M 65.17 15.82 39.02 -
+PiSSA 89.1M 67.01 18.13 40.85 -
+Bone 88.1M 66.94 18.85 41.76 -
 4.3 Effect of Rank $r$
@@
 As shown in the table, Bat, equipped with “Weight Guide”, achieves significant improvements in performance metrics compared to Bone.
-Table 4: Comparing Bat-row, Bat-col on math tasks
+Table 4: Comparing Bone, Bat on math tasks
 Model Strategy Trainable GSM8K Math
@@
 Bat 87.0M 49.36 8.88
-RWKV6-3B Bone 55.1M 41.73 6.52
+RWKV6-7B Bone 55.1M 41.73 6.52
 Bat 55.1M 42.76 6.60
@@
 RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
-Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.95}$ kt/s
+Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.99}$ kt/s
 Bat 55.1M 22978 MB 2.16 kt/s

### v6

@@
-Bone: Block-Affine Adaptation of Large Language Models
+DiSHA: Dimension-Sharding Adaptation with Fast Convergence and Fast Computation
 Abstract
-Low-Rank Adaptation (LoRA) has achieved remarkable training results by freezing the original weights and training only low-rank matrices, establishing itself as the predominant fine-tuning method for LLMs.
-Many LoRA variants have emerged, yet they lack a design tailored to the characteristics of LLM weights and fail to leverage the original weights effectively.
-To address the sparsity of LLM weights, and drawing inspiration from GQA and MQA, we propose Block-Affine Adaptation (Bone), a novel PEFT technique distinct from LoRA.
-By dividing the original weights into multiple subspaces that share a single matrix for weight updates, Bone simplifies the process by requiring the trainable matrix to be initialized to zero, eliminating the need for complex initialization as in some LoRA variants.
-Compared to LoRA, Bone significantly reduces memory usage and achieves faster computation.
-Evaluation of both NLU and NLG tasks demonstrates that Bone substantially outperforms LoRA and its variants.
-Inspired by Pissa, we propose a new theory called “Weight Guide” to better utilize the information embedded in the original weights.
-This approach extracts valuable information through a linear transformation of the original weight matrix using a trainable matrix.
-To validate the effectiveness of “Weight Guide” we combined it with Bone to create a new structure called Block-Affine Transformation (Bat), and ablation experiments confirmed the effectiveness of “Weight Guide”.
+Low-Rank Adaptation (LoRA) leverages the low intrinsic rank of weight updates in Large Language Models (LLMs), establishing a Parameter-Efficient Fine-Tuning (PEFT) paradigm.
+However, LoRA suffers from slow convergence.
+We introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to unlock lower intrinsic ranks and faster convergence by default.
+Within DiSHA’s design space, we propose Block Affine Adaptation (Bone), a computationally efficient structure that delivers both high performance and efficiency.
+While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation Adaptation (BAT), a nonlinear variant of DiSHA.
+BAT introduces nonlinearity by combining trainable matrices with original weight shards in a nonlinear manner, inducing nonlinearity in matrix updates without introducing additional parameters.
+Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both NLG and NLU tasks, with significantly improved computational efficiency.
+Further analysis demonstrates that BAT enhances model capabilities by leveraging its nonlinear design.
 1 Introduction
-Large models have been integrated into various industries, revolutionizing many traditional technologies Radford et al.
+The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  Radford et al.
 (2019); Raffel et al.
@@
 However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios.
-Full-scale fine-tuning of large models is computationally costly; for example, finetuning the LLaMA2-7B Touvron et al.
-(2023) model with bfloat16 Wang & Kanwar (2019) precision requires around 60GB of VRAM.
-In contrast, PEFT (Parameter-Efficient Fine-Tuning)Xu et al.
-(2023) techniques could reduce the VRAM requirement to fit into a 24GB VRAM GPU.
-As a result, numerous PEFT techniques and quantization methods have emerged to reduce the training costs of large models.
-LoRA (Low-Rank Adaptation) Hu et al.
-(2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after finetuning.
-LoRA significantly reduces memory usage by freezing the original weights $W$ and updating two low-rank matrices $A$ and $B$.
-Typically, either $A$ or $B$ is initialized to zero, ensuring that the initial state of LoRA is consistent with the pre-trained model, The figure  2 illustrates the structure visualization.
+Full-scale fine-tuning of large models is computationally costly.
+As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)Xu et al.
+(2023) techniques and quantization methods have emerged to reduce the training costs of large models.
+Low-Rank Adaptation(LoRA) Hu et al.
+(2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning.
+The figure  3 illustrates the structure visualization.
+32641282560612182430rankgradient normftfrozen ftLoRABigLoRAPiSSADiSHA
+Figure 1: the initial gradient norm of model training.
 However, extensive experiments Ding et al.
@@
 (2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning.
-This slow convergence is likely due to the small gradients caused by the zero initialization of either $A$ or $B$.
 To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al.
@@
 (2024).
-Despite their excellent performance, these LoRA variants inevitably introduce complexity into the fine-tuning process and reduce structural flexibility.
-For example, LoRA+ requires manual adjustment of different learning rates for $A$ and $B$; PISSA necessitates performing SVD decomposition on $W$ at the beginning of training, which can be time-consuming when the model parameters are large.
-As research on LoRA becomes increasingly saturated, it is essential to explore new directions for PEFT techniques and effectively leverage the unique characteristics of LLMs.
-In this work, we first designed Block-Affine Adaptation (Bone), which divides the original weights into multiple subspaces that share a single trainable matrix for updates, making it more efficient than LoRA’s two low-rank matrices.
-Secondly, to address the limitations of LoRA and fully utilize the information in the original weights, we propose a new theory, “Weight Guide”.
-“Weight Guide” enables feature extraction by applying a simple linear transformation to the original weights.
-Finally, to validate the effectiveness of “Weight Guide”, we combined Bone with “Weight Guide” to develop Block-Affine Transformation (Bat).
-Our extensive evaluation shows that Bone not only excels in both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks but also significantly outperforms LoRA and its variants.
-Additionally, Bone retains the advantages of LoRA, such as ease of use and no additional computational overhead for the model.
+These methods have all brought unexpected improvements.
+By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
+Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients.
+However, this also makes the initialization process significantly more complex, leading to increased initialization time overhead.
+To address the issue of complexity and explore the specific reasons for LoRA’s slow convergence, we conducted extensive experiments on initial gradient testing.
+Ultimately, we concluded that LoRA’s slow convergence is influenced not only by initialization but also by its inherent structural limitations.
+A simpler update structure may be more suitable for fine-tuning LLMs.
+In this work, we first propose a novel framework called Dimension-Sharding Adaptation (DiSHA) to design PEFT with lower ranks and faster convergence.
+In DiSHA, the pre-trained weights $W_{0}$ are frozen and divided into multiple shards during training, with all shard weights sharing a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters.
+The design space of DiSHA is vast, allowing pre-trained weights to be divided into dimensions of arbitrary sizes 2.
+Computational efficiencies vary among design choices.
+We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
+Of course, we also observed that certain subsets of the DiSHA design space exhibit limitations: all shard weight updates are colinear, which could potentially hinder model performance.
+To address this potential issue, we propose the Bat method.
+Without introducing additional trainable matrices, Bat uses the trainable matrix to interact with the original weight shards, transforming the linear updates between the shards into non-linear updates.
+This approach enhances the model’s capability while maintaining efficiency.
+Figure 2: partition of weight matrix $W_{0}$.
+Our extensive evaluation shows that Bone demonstrates excellent performance on both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks, significantly surpassing LoRA and its variants.
+Moreover, as shown in Figure  5, Bone also achieves a significantly faster convergence rate compared to PiSSA.
+This demonstrates the feasibility of the DiSHA framework.
+Furthermore, Bone’s unique design not only resolves the unification challenges of the DiSHA framework, allowing seamless adaptation across different LLM architectures such as LLaMA, RWKV, and Mistral but also enables it to significantly outperform LoRA in both computational efficiency and memory usage.
 In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board.
@@
 (2024)).
-As shown in Figure  1, Bone achieves the fastest convergence, and the results on the test set demonstrate its superior data fitting and generalization capabilities ( Table  2).
-As a completely new structure distinct from LoRA, Bone offers clear improvements in computational efficiency and memory savings, as detailed in table  7.
+Based on the comparative experiments between Bone and Bat, it is evident that the non-linear updates of weight shards enhance the model’s performance ceiling.
 Our contributions can be summarized as follows
-1.
-We propose a novel PEFT technique called Block-Affine Adaptation (Bone).
-Bone outperforms LoRA and its variants across various fine-tuning tasks.
-Additionally, Bone is more memory-efficient and computationally faster compared to LoRA.
-2.
-To effectively leverage the implicit information in the original weights, we propose the “Weight Guide” theory.
-By integrating this theory with Bone, we design Block-Affine Transformation (Bat), demonstrating the effectiveness of “Weight Guide”.
-(a) llama2-7b
-(b) Train Steps
-Figure 1: The left image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation.
-Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings.
-Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
-The left image is a demonstration of the training state.
+1. We propose a novel PEFT framework called DiSHA which significantly reduces the number of trainable parameters.
+2. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
+3.
+We identified a potential issue in the DiSHA framework: the updates of weight shards are colinear.
+To address this, we improved upon Bone and proposed Bat, which transforms weight shard updates into non-linear ones, thereby enhancing the model’s performance ceiling.
+4. Through extensive experiments, we validated the effectiveness of DiSHA, Bone, and Bat, achieving state-of-the-art (SOTA) performance.
 2 Related Works
@@
 (2024), and layer-freezing methods such as LISA.
-The adapter method does not require fine-tuning all the parameters of the pre-trained model.
-Instead, it introduces a small number of task-specific parameters to store knowledge related to that task, thereby reducing the computational demands of model fine-tuning.
-Prefix tuning is a lightweight fine-tuning method for generative tasks.
-It adds a continuous, task-specific vector sequence, called a prefix, to the input.
-Unlike prompts, prefixes are entirely composed of free parameters and do not correspond to actual tokens.
-Compared to traditional fine-tuning, prefix tuning only optimizes a set of prefixes related to specific tasks, without any change to the original model.
-Prompt tuning defines a unique prompt for each task, prepending it to the input data while freezing the pre-trained model during training.
-LoRA allows us to indirectly train certain dense layers in a neural network by optimizing low-rank decomposition matrices that adapt these layers while keeping the pre-trained weights unchanged.
-This approach significantly addresses the inference latency introduced by adapters and the fitting challenges of prefix tuning.
-In LoRA, the adapter matrices $A$ and $B$ are updated with the same learning rate, but using the same rate for both may not effectively learn the features.
+LoRA, as one of the most popular PEFT techniques, has demonstrated that the updates to the weights exhibit a low intrinsic rank during adaptation.
+In LoRA, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$.
+During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters.
+Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise.
+For $h=W_{0}x$, our modified forward pass yields:
+$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (1)
+Since A and B are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process.
+As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial gradients, among other approaches.
 LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency.
@@
 LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
+The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential.
+However, they also introduce more complex initialization steps and increase preprocessing time.
+Therefore, we need to design novel structures to facilitate the exploration of a lower intrinsic rank.
 3 Method
-3.1 Motivation
-The primary motivations are as follows:
-Firstly, current research on LoRA has reached a saturation point, underscoring the need for a new direction in Parameter-Efficient Fine-Tuning (PEFT) techniques.
-Secondly, existing PEFT methods lack a thorough exploration and effective utilization of the original model weights.
-3.2 Bone: Block-Affine update matrices
-To design a novel PEFT structure that rivals LoRA, we conducted an in-depth study of dense matrices.
-Current theories suggest that LLM weights exhibit sparse properties.
-For example, LoRA utilizes low-rank matrices to update dense matrices, while techniques like GQA Ainslie et al.
-(2023) and MQA Shazeer (2019) employ shared mechanisms to reduce parameter overhead with remarkable effectiveness.
-Inspired by these approaches, we hypothesize that dividing LLM weights into multiple subspaces allows different subspaces to share a single low-rank matrix for updates.
-As illustrated in the figure  2, we name this structure Block-Affine Adaptation (Bone).
-It is evident that the ways to partition LLM weights are diverse, as illustrated in the figure.
-Different partitioning methods influence the shape of the Bone matrix, the computation logic, and ultimately, the processing speed.
-In this work, we propose the Best Bone structure.
-For a pre-trained weight matrix $W\in\mathbb{R}^{d\times k}$, we update it using a single matrix: $W+\Delta W=W+expand(B)$, where $B\in\mathbb{R}^{r\times d}$, and the rank $r\ll k$.
-This design not only retains the simplicity and ease of use characteristic of LoRA but also achieves higher computational efficiency.
-The calculation is expressed as follows:
-$Y=Wx+\Delta Wx=Wx+sum(x)B$ (1)
-(a) Bone instruciton
-Figure 2: Bone requires training only a single low-rank matrix initialized to zero. Best Bone represents the optimal design structure of Bone. Subsequent experiments default to using Best Bone.
-3.3 “Weight Guide”
-It is well-known that LLMs have extremely large parameters, with complex interactions among their internal weights, but most LoRA variants utilize original weight information minimally, with only a few leveraging it during initialization to enhance performance.
-To increase the utilization of this information and promote internal interactions among weights, we propose a novel concept called “Weight Guide” This approach ensures that the trainable matrices are consistently guided and constrained by the original weights throughout every step of the training process.
-As illustrated in Figure  1(b), LoRA variants interact with the original weights only once during initialization to extract essential components.
-In contrast, the Bone structure is guided (constrained) by the original weights after each update, significantly increasing the new weights’ utilization of the original weights.
-As the training steps increase, the influence of the original weights becomes more pronounced.
-This is evident from the loss curves, which show that the differences between Bone and other methods grow larger in the later stages of training.
-So, how can we extract the critical information from the original weights?
-A straightforward approach is to apply a learnable matrix directly to the original weights using the Hadamard product.
-This method essentially assigns a weight to each element, allowing the learnable matrix to autonomously determine the importance of individual elements.
-Although this method is simple and straightforward, the elements within the weights remain isolated, making it impossible to learn interactions between them.
-Therefore, we replace the Hadamard product with matrix multiplication, essentially applying a simple linear transformation.
-This approach links the entire weight matrix space, providing a better chance to discover the optimal solution.
-The corresponding formula is shown below:
-$\Delta W_{n,n}$ $=W_{n,n}\odot bone_{n,n}$ (2)
-$\Delta W_{n,n}$ $=W_{n,n}\otimes bone_{n,n}$
+3.1 DiSHA: Dimension-Sharding Adaptation
+Hypothesis Based on the conclusions of the experimental analysis of the gradient norm  9.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower intrinsic ranks and faster convergence by default.
+For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W=W_{0}+\Delta W=W_{0}+expand(D)$, where $D\in\mathbb{R}^{r_{1}\times r_{2}},expand(D)\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$.
+During training, $W_{0}$ is frozen and does not receive gradient updates, while $D$ is a trainable parameter initialized to 0.
+Note both $W_{0}$ and $\Delta W=expand(D)$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise.
+For $y=W_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
+$y=W_{0}x+\Delta Wx=W_{0}x+expand(D)x$ (2)
+Here, we explain the concept of $expand(D)$.
+As shown in the DiSHA structure in Figure  3, we divide $W_{0}$ into multiple shards along the dimension of $D$.
+The $D$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $W_{0}$.
+Finally, the updated shards are summed to update $W$.
+3.2 DiSHA design space
+In the previous section, we introduced the design principles of DiSHA and defined its trainable matrix $D\in\mathbb{R}^{r_{1}\times r_{2}}$.
+With DiSHA, you can customize the size of the trainable matrix $D$ based on the dimensions of the pre-trained weights.
+Since all shards of the pre-trained weights $W_{0}$ share a single $D$ for updates, as long as the expanded $expand(D)$ matches the dimensions of $W_{0}$, you can freely choose how to divide $W_{0}$ into shards.
+This flexibility makes the design space of DiSHA immense 2.
+3.3 Bone: Block-Affine efficient update
+In the previous section, we introduced the DiSHA framework, a method that significantly reduces the number of trainable parameters in the model.
+Due to the vast design space of the DiSHA framework, it is difficult for users to accurately determine suitable values for training.
+A good setting of $D\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks.
+For example, with the rise of technologies like GQA Ainslie et al.
+(2023) and MQA Shazeer (2019), the matrix dimensions of self-attention modules ($W_{k},W_{v}$) vary greatly among different LLM models.
+At the same time, computational efficiency and resource utilization also vary depending on the design space choices within the DiSHA framework.
+To address these limitations, we proposed Block Affine Adaptation (Bone) based on the DiSHA framework.
+Inspired by LoRA, we modified $D\in\mathbb{R}^{r_{1}\times r_{2}}$ to $D\in\mathbb{R}^{r\times d}$ or $D\in\mathbb{R}^{r\times k}$.
+This allows the trainable parameter count to be controlled by a single parameter $r$, making it easy to adapt to any neural network.
+Additionally, we observed that $expand(D)x$ not only increases memory consumption but also reduces computational efficiency.
+From our analysis, when $D\in\mathbb{R}^{r\times d}$, $x\in\mathbb{R}^{b\times l\times k}$, $expand(D)\in\mathbb{R}^{d\times k}$, the operation $expand(D)x$ essentially performs a summation along the $k$-dimension, where $expand(D)$ involves replicating the necessary number of shards.
+Clearly, we can first $sum(x)$ along the $k$dimension and then compute it directly with $D$, where $sum(x)\in\mathbb{R}^{b\times l\times r}$.
+This approach not only reduces memory consumption but also improves computational efficiency.
+The update for $W$ is shown in the figure  3, and the Bone calculation formula is as follows:
+$y=W_{0}x+\Delta Wx=W_{0}x+sum(x)D$ (3)
+Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
 3.4 Bat: Block-Affine Transformation
-In the previous section, we introduced Bone, which is both simple and efficient.
-However, applying the same updates to different subspaces of the weight matrix $W\in\mathbb{R}^{d\times k}$ is clearly counterintuitive.
-To allow for more flexibility in the update directions of different subspaces, we combine Bone with “Weight Guide”, proposing a new method called Block-Affine Transformation (Bat).
-To balance computational efficiency, we divide the original weights into smaller subspaces compared to Bone.
-Here, we name the trainable matrix $bone\in\mathbb{R}^{r\times d}$, where its rank $r$ is replaced with block size $b$.
-The formula for Bone-col is as follows:
-$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (3)
-$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
-$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\otimes bone_{d/b,b,b}+bone_{d/b,b,b})$
-Similar to Bone, Bat can also adopt various grouping methods, as illustrated in the figure 3.
+We conducted extensive experiments on both NLU and NLG tasks to validate the effectiveness of Bone.
+It outperforms many LoRA variants and surpasses LoRA in terms of memory consumption and computational efficiency.
+However, Bone still has a potential issue, which stems from the shared update mechanism of the DiSHA framework.
+Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be colinear.
+This restricts weight updates and, consequently, the model’s capabilities.
+To address the issue of linear correlation among shard updates, our initial idea was to use a trainable coefficient matrix to control the updates of different shards.
+However, this approach would increase additional parameters.
+To avoid introducing additional parameters, we applied a linear transformation between the original weight $W_{0}$ shards and the trainable matrix, ensuring that the updates of different shards are nonlinearly related.
+Based on this, we proposed the Block Affine Transformation Adaptation (Bat) method.
+It extends Bone by further dividing both the trainable matrix and $W_{0}$ into smaller shards, where each shard interacts with the trainable matrix $D$ through a linear transformation, achieving nonlinear independence among shard updates.
+The specific formula is as follows:
+$W_{k/r,d/r,r,r}$ $=Reshape(W_{d,k})$ (4)
+$bone_{d/r,r,r}$ $=Reshape(bone_{r,d})$
+$\Delta W_{d,k}$ $=Reshape(W_{k/r,d/r,r,r}bone_{d/r,r,r}+bone_{d/r,r,r})$
+It is evident that Bat can be defined in many other ways, as illustrated in the figure 4.
 In Section 6, we compared the performance of different partitioning strategies for the structure.
 The formula for Bone-row is as follows:
-$W_{d/b,k/b,b,b}$ $=Reshape(W_{d,k})$ (4)
-$bone_{k/b,b,b}$ $=Reshape(bone_{b,k})$
-$\Delta W_{d,k}$ $=Reshape(W_{d/b,k/b,b,b}\otimes bone_{k/b,b,b}+bone_{k/b,b,b})$
-(a) Bone-both
-(b) Bone-col
-(c) Bone-row
-[…]
+$W_{d/r,k/r,r,r}$ $=Reshape(W_{d,k})$ (5)
+$bone_{k/r,r,r}$ $=Reshape(bone_{r,k})$
+$\Delta W_{d,k}$ $=Reshape(W_{d/r,k/r,r,r}bone_{k/r,r,r}+bone_{k/r,r,r})$
+Figure 4: Diagram of Bat Structures in Different Design Spaces.
 4 Experiments
@@
 We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
+Figure 5: The image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation.
+Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings.
+Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
 Implementation Details
@@
 Result
-The results, as shown in Table  2 and Figure  1(a), demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
+The results, as shown in Table  2 and Figure  5, demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
 Bone demonstrates outstanding performance across three different tasks.
@@
 4.4 Bone vs Bat
-To validate the effectiveness of the “Weight Guide”, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K.
-As shown in the table, Bat, equipped with “Weight Guide”, achieves significant improvements in performance metrics compared to Bone.
 Table 4: Comparing Bone, Bat on math tasks
@@
 Bat 55.1M 42.76 6.60
+In this section, we have verified the effectiveness of non-linear updates by comparing the performance of Bone and Bat, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K.
+As shown in the table, with the same number of trainable parameters, the metrics of non-linear updates are better, successfully breaking the original limits.
 4.5 Ablation experiments on different grouping methods for Bat
@@
 This allows for a fairer comparison between Bat-row and Bat-col.
+Table 5: Comparing Bat-row, Bat-col on math tasks
+Model Strategy Trainable GSM8K Math
+Llama2-7B Bat-row 72.8M 45.76 7.82
+Bat-col 87.0M 49.36 8.88
+RWKV6-3B Bat-row 55.1M 25.93 3.12
+Bat-col 55.1M 25.25 3.09
 The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well.
@@
 (2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
-Table 5: Comparing Bat-row, Bat-col on math tasks
-Model Strategy Trainable GSM8K Math
-Llama2-7B Bat-row 72.8M 45.76 7.82
-Bat-col 87.0M 49.36 8.88
-RWKV6-3B Bat-row 55.1M 25.93 3.12
-Bat-col 55.1M 25.25 3.09
-4.6 Ablation experiments on the specific implementation of “Weight Guide”
-Evaluating the best computation method for promoting internal weight feature fusion on RWKV6-3B: As seen in the table  6, Bone-Hadamard is significantly weaker compared to matrix multiplication.
-The formula for Bone-Hadamard is as follows:
-$W_{k/b,d/b,b,b}$ $=Reshape(W_{d,k})$ (5)
-$bone_{d/b,b,b}$ $=Reshape(bone_{b,d})$
-$\Delta W_{d,k}$ $=Reshape(W_{k/b,d/b,b,b}\odot bone_{d/b,b,b}+bone_{d/b,b,b})$
-Table 6: Block-Affine on math tasks
-Model Strategy Trainable GSM8K Math
-RWKV6-3B Bat 55.1M 25.93 3.32
-Bat-Hadamard 55.1M 22.44 2.67
 5 Resource and efficiency
-Table  7 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU.
+Table  6 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU.
 The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512.
@@
 At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
-Therefore, a key focus of our future work will be improving the Bat operator to enhance token throughput and reduce memory usage.
 Table 6: Resource and efficiency
@@
 6 Conclusion
-This work, inspired by GQA and MQA, leverages the sparsity of LLM weights to design the Block-Affine Adaptation (Bone) structure.
-In Bone, the original weights are divided into multiple subspaces, all of which share a single low-rank matrix initialized to zero for updates.
-Extensive experiments demonstrate that Bone consistently outperforms LoRA and its variants across various tasks, while also offering superior computational efficiency.
-To break the limitations of LoRA and effectively utilize the information from the original weights, we propose the “Weight Guide” theory, which enables significant improvements through simple linear transformations.
-By integrating these innovations, we introduce a new structure called Block-Affine Transformation (Bat).
-Bat not only validates the effectiveness of “Weight Guide” but also addresses the limitation of Bone, where identical updates are applied to all subspaces.
-Experimental results show that Bat surpasses Bone on multiple tasks, although with a trade-off in computational efficiency.
-Bone brings new possibilities to existing LLM PEFT techniques.
-Instead of focusing solely on optimizing LoRA, we should shift our attention to innovative PEFT methods that are better suited to the architecture of LLMs.
+In this work, we proposed the DiSHA framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix.
+This approach significantly reduces resource overhead and opens up new directions for PEFT techniques.
+Building on the DiSHA framework, we introduced Bone, a broad and efficient method comparable to LoRA.
+Extensive experiments demonstrated the superiority of Bone, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
+To address the potential issue of linear correlation in shard updates within Bone and the DiSHA framework, we proposed Bat.
+By applying a linear transformation between the trainable matrix and the pre-trained weights, Bat ensures nonlinear independence in the updates of different shards, successfully breaking through the structural limitations of Bone
 7 Future work
-Bone and “Weight Guide” are merely a starting point, offering a foundation for researchers to explore additional branches inspired by Bat. There are many directions for future work.
-1. Can Bone adapt to multimodal tasks and other complex scenarios?
-2. Can “Weight Guide” be implemented using alternative computation methods?
+Due to the limitations of the experimental equipment, we did not report any results from the full-scale training required for this work.
+However, through extensive evaluations and comparisons with LoRA and its variants, we validated the effectiveness of DiSHA, Bone, and Bat.
+The outstanding performance demonstrated in LLMs by this work gives us great confidence.
+We believe that DiSHA, Bone, and Bat can be applied to any multimodal task.
+Of course, this work also has some limitations that need to be addressed.
+Although Bat successfully overcomes the limitations of Bone, it also leads to reduced computational efficiency and increased resource consumption.
+Therefore, in the future, we plan to develop specialized optimized operators to address this issue.
 We welcome the community to provide additional suggestions and conduct further tests.
+8 Acknowledgments
+I would like to express my heartfelt gratitude to all the seniors who provided valuable feedback on my work.
+Their insightful suggestions have greatly enhanced the quality of this paper.
+I also wish to thank RWKVOS for their support.
+Finally, I sincerely appreciate the meticulous guidance provided by the anonymous online mentor, Professor Cormen, whose help has been indispensable.

### v7

@@
-DiSHA: Dimension-Sharding Adaptation with Fast Convergence and Fast Computation
+DiSHA: Dimension-Sharding Adaptation of Large Language Models with Fast Convergence and Fast Computation
 Abstract
-Low-Rank Adaptation (LoRA) leverages the low intrinsic rank of weight updates in Large Language Models (LLMs), establishing a Parameter-Efficient Fine-Tuning (PEFT) paradigm.
-However, LoRA suffers from slow convergence.
-We introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to unlock lower intrinsic ranks and faster convergence by default.
-Within DiSHA’s design space, we propose Block Affine Adaptation (Bone), a computationally efficient structure that delivers both high performance and efficiency.
-While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation Adaptation (BAT), a nonlinear variant of DiSHA.
+Low-Rank Adaptation (LoRA), a prominent technique within the framework of Parameter-Efficient Fine-Tuning (PEFT), efficiently reduces the computational burden associated with adapting Large Language Models (LLMs) to downstream tasks, thereby enabling resource-constrained fine-tuning.
+However, existing researches have shown that LoRA suffers from slow convergence.
+To address this limitation, we introduce Dimension-Sharding Adaptation (DiSHA), which expands the PEFT design space to even fewer trainable parameters and faster convergence.
+Within DiSHA’s design space, we propose Block Affine Efficient Computation (Bone), a computationally efficient structure that delivers both high performance and efficiency.
+While certain DiSHA configurations may result in colinear updates to weight shards, we address this with Block Affine Transformation (Bat), a nonlinear variant of DiSHA.
 Bat introduces nonlinearity by combining trainable matrices with original weight shards in a nonlinear manner, inducing nonlinearity in matrix updates without introducing additional parameters.
-Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both NLG and NLU tasks, with significantly improved computational efficiency.
+Empirical results show that Bone, under the DiSHA framework, consistently outperforms LoRA variants in both Natural Language Understanding and Natural Language Generation tasks, with significantly improved computational efficiency.
 Further analysis demonstrates that BAT enhances model capabilities by leveraging its nonlinear design.
+[Uncaptioned image]https://github.com/huggingface/peft
+Machine Learning, ICML
 1 Introduction
-The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  Radford et al.
-(2019); Raffel et al.
-(2020).
+The emergence of Large Language Models (LLMs) has fundamentally transformed many traditional technologies  (Radford et al., 2019; Raffel et al., 2020).
 However, general-purpose large models often struggle to meet the needs of all downstream tasks, making it necessary to fine-tune base models for specific scenarios.
 Full-scale fine-tuning of large models is computationally costly.
-As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)Xu et al.
-(2023) techniques and quantization methods have emerged to reduce the training costs of large models.
-Low-Rank Adaptation(LoRA) Hu et al.
-(2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning.
+As a result, numerous Parameter-Efficient Fine-Tuning (PEFT)(Xu et al., 2023) techniques and quantization methods have emerged to reduce the training costs of large models.
+Low-Rank Adaptation(LoRA) (Hu et al., 2021) has become one of the most popular PEFT methods due to its small tunable parameter size, its effectiveness, and the possibility of zero inference overhead after fine-tuning.
 The figure  3 illustrates the structure visualization.
@@
 Figure 1: the initial gradient norm of model training.
-However, extensive experiments Ding et al.
-(2023); Liu et al.
-(2024b); Biderman et al.
-(2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning.
-To address this issue, researchers have proposed several LoRA variants, such as LoRA+ Hayou et al.
-(2024), PISSA Meng et al.
-(2024), and LoRA-GAWang et al.
-(2024).
+However, extensive experiments (Ding et al., 2023; Liu et al., 2024b; Biderman et al., 2024) have shown that LoRA’s convergence is significantly slower compared to full fine-tuning.
+To address this issue, researchers have proposed several LoRA variants, such as LoRA+ (Hayou et al., 2024), PISSA (Meng et al., 2024), and LoRA-GA(Wang et al., 2024).
 These methods have all brought unexpected improvements.
 By adopting different initialization strategies to influence the model’s training gradients, they have accelerated LoRA’s convergence speed.
+Figure 2: partition of weight matrix $\mathbf{W_{0}}$.
 Different initializations of LoRA variants accelerate convergence essentially by increasing the initial gradients during training or aligning them with the full-scale training gradients.
@@
 A simpler update structure may be more suitable for fine-tuning LLMs.
-In this work, we first propose a novel framework called Dimension-Sharding Adaptation (DiSHA) to design PEFT with lower ranks and faster convergence.
-In DiSHA, the pre-trained weights $W_{0}$ are frozen and divided into multiple shards during training, with all shard weights sharing a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters.
-The design space of DiSHA is vast, allowing pre-trained weights to be divided into dimensions of arbitrary sizes 2.
-Computational efficiencies vary among design choices.
-We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
-Of course, we also observed that certain subsets of the DiSHA design space exhibit limitations: all shard weight updates are colinear, which could potentially hinder model performance.
-To address this potential issue, we propose the Bat method.
-Without introducing additional trainable matrices, Bat uses the trainable matrix to interact with the original weight shards, transforming the linear updates between the shards into non-linear updates.
-This approach enhances the model’s capability while maintaining efficiency.
-Figure 2: partition of weight matrix $W_{0}$.
+Based on the aforementioned conjecture, we first propose a framework with a simpler update mechanism, called Dimension-Sharding Adaptation (DiSHA), designed to achieve larger initial gradients with fewer trainable parameters, thereby accelerating the convergence process.
+In DiSHA, the pre-trained weight matrix $\mathbf{W_{0}}$ is frozen and partitioned into multiple shard matrices during training, and all shard matrices share a trainable matrix of the same size as a shard to perform updates, significantly reducing the number of trainable parameters.
+The design space of DiSHA is extensive, permitting the division of pre-trained weights matrix into dimensions of arbitrary magnitudes, as illustrated in Figure 2.
+Computational efficiencies vary among design choices, so we proposed an efficient and unified structure named Block Affine Efficient Computation (Bone) within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
+As the experiments progressed, we discovered that when fine-tuning with Bone, the updates between different shards within the same matrix are collinear; therefore, to introduce nonlinear updates, we proposed the Block Affine Transformation (Bat) method.
+Inspired by methods like PiSSA that leverage pre-trained weight matrix information, the updates between different shard matrices are no longer directly added to the trainable matrix.
+Instead, they need to be linearly transformed with the corresponding shard matrix before being summed.
+This approach not only leverages the information from the pre-trained weight matrices but also ensures that the updates between different shard matrices become nonlinear.
 Our extensive evaluation shows that Bone demonstrates excellent performance on both Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks, significantly surpassing LoRA and its variants.
@@
 In NLG tasks, Bone demonstrates superior performance, with evaluation metrics surpassing even the strong LoRA variant Pissa across the board.
-To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 Xu et al.
-(2023), RWKV6 Peng et al.
-(2024)).
+To verify the feasibility of Bone, we conducted experiments on two different LLM architectures (LLaMA2 (Xu et al., 2023), RWKV6 (Peng et al., 2024)).
 Based on the comparative experiments between Bone and Bat, it is evident that the non-linear updates of weight shards enhance the model’s performance ceiling.
@@
 1. We propose a novel PEFT framework called DiSHA which significantly reduces the number of trainable parameters.
-2. We proposed an efficient and user-friendly structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
+2. We proposed an efficient and adaptable structure named Bone within a subset of the DiSHA design space and validated its effectiveness through extensive experiments.
 3.
-We identified a potential issue in the DiSHA framework: the updates of weight shards are colinear.
-To address this, we improved upon Bone and proposed Bat, which transforms weight shard updates into non-linear ones, thereby enhancing the model’s performance ceiling.
+We propose Block Affine Transformation (Bat), which introduces shard-specific nonlinearity by integrating pre-trained weight information, without adding new parameters: A breakthrough to overcome the collinear update bottleneck inherent in both Bone.
 4. Through extensive experiments, we validated the effectiveness of DiSHA, Bone, and Bat, achieving state-of-the-art (SOTA) performance.
 2 Related Works
-The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuningHoulsby et al.
-(2019); He et al.
-(2022); Wang et al.
-(2022); Pfeiffer et al.
-(2020), prefix tuningLiu et al.
-(2023); Li & Liang (2021), prompt tuningBrown (2020); Liu et al.
-(2023); Lester et al.
-(2021); Razdaibiedina et al.
-(2023); Li & Liang (2021), LoRAHu et al.
-(2021); Meng et al.
-(2024); Wang et al.
-(2024); Si et al.
-(2024), and layer-freezing methods such as LISA.
+The PEFT (Parameter-Efficient Fine-Tuning) techniques are diverse and include approaches like adapter tuning(Houlsby et al., 2019; He et al., 2022; Wang et al., 2022; Pfeiffer et al., 2020), prefix tuning(Liu et al., 2023; Li & Liang, 2021), prompt tuning(Brown, 2020; Liu et al., 2023; Lester et al., 2021; Razdaibiedina et al., 2023; Li & Liang, 2021), LoRA(Hu et al., 2021; Meng et al., 2024; Wang et al., 2024; Si et al., 2024), and layer-freezing methods such as LISA.
 LoRA, as one of the most popular PEFT techniques, has demonstrated that the updates to the weights exhibit a low intrinsic rank during adaptation.
-In LoRA, for a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W_{0}+\Delta W=W_{0}+BA$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$.
-During training, $W_{0}$ is frozen and does not receive gradient updates, while $A$ and $B$ contain trainable parameters.
-Note both $W_{0}$ and $\Delta W=BA$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise.
-For $h=W_{0}x$, our modified forward pass yields:
-$h=W_{0}x+\Delta Wx=W_{0}x+BAx$ (1)
+In LoRA, for a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $\mathbf{W}_{0}+\mathbf{\Delta W}=\mathbf{W}_{0}+\mathbf{BA}$, where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, and the rank $r\ll\min(d,k)$.
+During training, $\mathbf{W}_{0}$ is frozen and does not receive gradient updates, while $\mathbf{A}$ and $\mathbf{B}$ contain trainable parameters.
+Note both $\mathbf{W}_{0}$ and $\mathbf{\Delta W}=\mathbf{BA}$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise.
+For $h=\mathbf{W}_{0}x$, our modified forward pass yields:
+$h=\mathbf{W}_{0}x+\mathbf{\Delta W}x=\mathbf{W}_{0}x+\mathbf{BA}x$ (1)
 Since A and B are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process.
-As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial gradients, among other approaches.
-LoRA+ extended this method by introducing independent learning rates for matrices $A$ and $B$ with a fixed ratio, improving the method’s efficiency.
-The DoRA Liu et al.
-(2024b) method combines weight decomposition to achieve learning capabilities similar to full fine-tuning without sacrificing LoRA’s inference efficiency.
+As a result, many LoRA variants have emerged, aiming to enhance LoRA’s capabilities by modifying the initialization or initial learning rat, among other approaches.
+2.1 Adjusting the initial learning rate of LoRA
+LoRA+ extended this method by introducing independent learning rates for matrices $\mathbf{A}$ and $\mathbf{B}$ with a fixed ratio, improving the method’s efficiency.
+DoRA (Liu et al., 2024b) decomposes the weight matrix into two parts: magnitude and direction, which are optimized separately.
+This approach allows for more precise control over the learning rate, making LoRA updates closer to the effect of full fine-tuning.
+2.2 Improving the Initialization of LoRA
 PiSSA optimizes the compact parameter space by representing the matrices in the model as the product of two trainable matrices, augmented with a residual matrix for error correction.
 Using Singular Value Decomposition (SVD), PiSSA initializes the dominant singular values and vectors to train these matrices, while keeping the residual matrix static during fine-tuning.
-OLoRA Büyükakyüz (2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal.
+OLoRA (Büyükakyüz, 2024) leverages QR decomposition to initialize the adaptation matrices during the fine-tuning process, ensuring that these matrices are orthogonal.
 This orthogonal initialization helps maintain the stability of the parameter space during optimization.
-LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $A$ and $B$ by computing the initial gradient, thereby closely approximating full fine-tuning.
-The improvements brought by these LoRA variants validate that the updates to the weights exhibit a low intrinsic rank during adaptation and hold greater potential.
-However, they also introduce more complex initialization steps and increase preprocessing time.
-Therefore, we need to design novel structures to facilitate the exploration of a lower intrinsic rank.
+LoRA-GA and PiSSA are similar in form, but they differ in that LoRA-GA initializes $\mathbf{A}$ and $\mathbf{B}$ by computing the initial gradient, thereby closely approximating full fine-tuning.
+Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
 3 Method
+3.1 Motivation: Issues with LoRA
+We have identified two main limitations of LoRA and its variants: Firstly, LoRA with simple initialization suffers from slow convergence.
+Secondly, while LoRA variants accelerate convergence, they also introduce many more complex operations.
+Specifically, methods like PiSSA and OLoRA, which set the initialization by decomposing the pre-trained weights, increase the initialization time.
+Moreover, the initialization time grows as the model’s parameter size increases.
 3.2 DiSHA: Dimension-Sharding Adaptation
-Hypothesis Based on the conclusions of the experimental analysis of the gradient norm  9.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower intrinsic ranks and faster convergence by default.
-For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $W=W_{0}+\Delta W=W_{0}+expand(D)$, where $D\in\mathbb{R}^{r_{1}\times r_{2}},expand(D)\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$.
-During training, $W_{0}$ is frozen and does not receive gradient updates, while $D$ is a trainable parameter initialized to 0.
-Note both $W_{0}$ and $\Delta W=expand(D)$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise.
-For $y=W_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
-$y=W_{0}x+\Delta Wx=W_{0}x+expand(D)x$ (2)
-Here, we explain the concept of $expand(D)$.
-As shown in the DiSHA structure in Figure  3, we divide $W_{0}$ into multiple shards along the dimension of $D$.
-The $D$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $W_{0}$.
-Finally, the updated shards are summed to update $W$.
+To address these limitations and based on the conclusions from the gradient norm experiments A.1, we designed a novel framework called Dimension-Sharding Adaptation (DiSHA) to unlock lower trainable parameters and faster convergence by default.
+For a pre-trained weight matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, we constrain its update by representing the latter with a low-rank decomposition $\mathbf{W}=\mathbf{W}_{0}+\mathbf{\Delta W}=\mathbf{W}_{0}+\operatorname{expand}(\mathbf{D})$, where $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}},\operatorname{expand}(\mathbf{D})\in\mathbb{R}^{d\times k}$, and the rank $(r_{1},r_{2})\ll\min(d,k)$.
+During training, $\mathbf{W}_{0}$ is frozen and does not receive gradient updates, while the trainable matrix $\mathbf{D}$ is initialized to zero, ensuring $\operatorname{expand}(\mathbf{D})$ initially contributes no update to $\mathbf{W}_{0}$.
+Note both $\mathbf{W}_{0}$ and $\mathbf{\Delta W}=\operatorname{expand}(\mathbf{D})$ are multiplied with the same input, and their respective output vectors are summed coordinate-wise.
+For $y=\mathbf{W}_{0}x$, where $x\in\mathbb{R}^{b\times l\times k}$, $y\in\mathbb{R}^{b\times l\times d}$, our modified forward pass yields:
+$y=\mathbf{W}_{0}x+\mathbf{\Delta W}x=\mathbf{W}_{0}x+\operatorname{expand}(\mathbf{D})x$ (2)
+Here, we explain the concept of $\operatorname{expand}(\mathbf{D})$.
+As shown in the DiSHA structure in Figure  3, we partition $\mathbf{W}_{0}$ into multiple shards along the dimension of $\mathbf{D}$.
+The $\mathbf{D}$ matrix is then replicated according to the number of shards, and its dimensions are restored to match $\mathbf{W}_{0}$.
+Finally, the updated shards are summed to update $\mathbf{W}_{0}$.
 3.3 DiSHA Design Space
-In the previous section, we introduced the design principles of DiSHA and defined its trainable matrix $D\in\mathbb{R}^{r_{1}\times r_{2}}$.
-With DiSHA, you can customize the size of the trainable matrix $D$ based on the dimensions of the pre-trained weights.
-Since all shards of the pre-trained weights $W_{0}$ share a single $D$ for updates, as long as the expanded $expand(D)$ matches the dimensions of $W_{0}$, you can freely choose how to divide $W_{0}$ into shards.
-This flexibility makes the design space of DiSHA immense 2.
-3.3 Bone: Block-Affine efficient update
+The design space of DiSHA is determined by its partitioned low-rank update mechanism.
+Given a pre-trained weight $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, the trainable matrix $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ governs updates through two configurable degrees of freedom:
+1. Rank Selection: The ranks $r_{1}$ and $r_{2}$ independently control the parameterization complexity. For example:
+• A wide matrix ($d\ll k$) benefits from $r_{1}\ll r_{2}$ to prioritize input-dimension compactness.
+• A tall matrix ($d\gg k$) favors $r_{1}\gg r_{2}$ to minimize memory overhead.
+2.
+Partitioning Strategy: The output dimension $d$ can be partitioned into $N$ shards with sizes $\{s_{1},...,s_{N}\}$, where $\sum_{i=1}^{N}s_{i}=d$.
+Each shard’s update is scaled by a row of $\mathbf{D}$:
+$\operatorname{expand}(\mathbf{D})=\bigoplus_{i=1}^{N}\left(\mathbf{1}_{s_{i}}\otimes\mathbf{D}_{i}\right)\in\mathbb{R}^{d\times k}$
+Here, $\mathbf{D}_{i}\in\mathbb{R}^{1\times r_{2}}$ is replicated $s_{i}$ times vertically.
+As shown in Table 2, DiSHA can set fewer trainable parameters while supporting non-uniform partitioning (e.g., grouping attention heads in transformers). This allows for a wide design space in DiSHA.
+3.4 Bone: Block Affine Efficient Computation
 In the previous section, we introduced the DiSHA framework, a method that significantly reduces the number of trainable parameters in the model.
 Due to the vast design space of the DiSHA framework, it is difficult for users to accurately determine suitable values for training.
-A good setting of $D\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks.
-For example, with the rise of technologies like GQA Ainslie et al.
-(2023) and MQA Shazeer (2019), the matrix dimensions of self-attention modules ($W_{k},W_{v}$) vary greatly among different LLM models.
-At the same time, computational efficiency and resource utilization also vary depending on the design space choices within the DiSHA framework.
-To address these limitations, we proposed Block Affine Adaptation (Bone) based on the DiSHA framework.
-Inspired by LoRA, we modified $D\in\mathbb{R}^{r_{1}\times r_{2}}$ to $D\in\mathbb{R}^{r\times d}$ or $D\in\mathbb{R}^{r\times k}$.
-This allows the trainable parameter count to be controlled by a single parameter $r$, making it easy to adapt to any neural network.
-Additionally, we observed that $expand(D)x$ not only increases memory consumption but also reduces computational efficiency.
-From our analysis, when $D\in\mathbb{R}^{r\times d}$, $x\in\mathbb{R}^{b\times l\times k}$, $expand(D)\in\mathbb{R}^{d\times k}$, the operation $expand(D)x$ essentially performs a summation along the $k$-dimension, where $expand(D)$ involves replicating the necessary number of shards.
-Clearly, we can first $sum(x)$ along the $k$dimension and then compute it directly with $D$, where $sum(x)\in\mathbb{R}^{b\times l\times r}$.
-This approach not only reduces memory consumption but also improves computational efficiency.
-The update for $W$ is shown in the figure  3, and the Bone calculation formula is as follows:
-$y=W_{0}x+\Delta Wx=W_{0}x+sum(x)D$ (3)
-Figure 3: Diagram of LoRA, DiSHA, and Bone Architectures.
-3.4 Bat: Block-Affine Transformation
+A good setting of $\mathbf{D}\in\mathbb{R}^{r_{1}\times r_{2}}$ does not generalize well across all neural networks.
+For example, with the rise of technologies like GQA (Ainslie et al., 2023) and MQA (Shazeer, 2019), the matrix dimensions of self-attention modules ($\mathbf{W}_{k},\mathbf{W}_{v}$) vary greatly among different LLM models.
+To enhance the adaptability and efficiency of DiSHA, we proposed Block Affine Efficient Computation (Bone), which reduces the memory footprint of intermediate parameters while enhancing computational efficiency through an innovative input aggregation mechanism.
+Core Innovation: Blockwise Input Aggregation Bone redefines the trainable matrix as $\mathbf{D}\in\mathbb{R}^{r\times d}$ with a single rank parameter $r$, enabling unified adaptation across heterogeneous layers.
+The key insight is to decouple the input dimension $k$ from the parameterization by introducing blockwise summation:
+1. Input Partitioning: For input $x\in\mathbb{R}^{b\times l\times k}$, partition the $k$-dimension into $r$ contiguous blocks of size $g=\lfloor k/r\rfloor$:
+$x\rightarrow[x^{(1)},\dots,x^{(r)}],\quad x^{(i)}\in\mathbb{R}^{b\times l\times g}$
+2. Blockwise Summation: Sum each partitioned block along the $k$-dimension:
+$\mathbf{S}=\left[\sum_{j=1}^{g}x^{(1)}_{:,:,j},\dots,\sum_{j=1}^{g}x^{(r)}_{:,:,j}\right]\in\mathbb{R}^{b\times l\times r}$
+3. Affine Projection: Compute the low-rank update via matrix multiplication with $\mathbf{D}$:
+$\mathbf{\Delta W}x=\mathbf{D}^{\top}\mathbf{S}\quad(\mathbf{D}\in\mathbb{R}^{r\times d},\mathbf{S}\in\mathbb{R}^{b\times l\times r})$
+The final forward pass becomes:
+$y=\mathbf{W}_{0}x+\mathbf{D}^{\top}\mathbf{S}$
+where $\mathbf{D}^{\top}\mathbf{S}\in\mathbb{R}^{b\times l\times d}$ matches the output dimension of $\mathbf{W}_{0}x$.
+Mathematical Equivalence to DiSHA Bone’s blockwise summation is equivalent to DiSHA’s $\operatorname{expand}(\mathbf{D})$ operation under structured sparsity.
+Let $\operatorname{expand}(\mathbf{D})\in\mathbb{R}^{d\times k}$ be constructed by tiling each row of $\mathbf{D}$ over $g$ columns:
+$\operatorname{expand}(\mathbf{D})_{i,j}=\mathbf{D}_{i,\lceil j/g\rceil}$
+Then:
+$\operatorname{expand}(\mathbf{D})x=\sum_{i=1}^{r}\mathbf{D}_{i}\sum_{j=1}^{g}x^{(i)}_{:,:,j}=\mathbf{D}^{\top}\mathbf{S}$
+This equivalence allows Bone to avoid explicitly materializing $\operatorname{expand}(\mathbf{D})$, reducing memory from $O(dk)$ to $O(dr)$, and reducing flops from $O(bldk)$ to $O(blr(d+\frac{k}{r}))$
+Figure 4: Diagram of Bat Structures in Different Design Spaces.
+3.5 Bat: Block Affine Transformation
 We conducted extensive experiments on both NLU and NLG tasks to validate the effectiveness of Bone.
 It outperforms many LoRA variants and surpasses LoRA in terms of memory consumption and computational efficiency.
-However, Bone still has a potential issue, which stems from the shared update mechanism of the DiSHA framework.
-Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be colinear.
-This restricts weight updates and, consequently, the model’s capabilities.
+However, we found that Bone results in the updates between different shards within the same matrix being collinear.
+Specifically, all shards in the weights use the same trainable matrix for updates, causing the updates of all shards to be collinear, restricting the model’s expressive power.
 To address the issue of linear correlation among shard updates, our initial idea was to use a trainable coefficient matrix to control the updates of different shards.
 However, this approach would increase additional parameters.
-To avoid introducing additional parameters, we applied a linear transformation between the original weight $W_{0}$ shards and the trainable matrix, ensuring that the updates of different shards are nonlinearly related.
-Based on this, we proposed the Block Affine Transformation Adaptation (Bat) method.
-It extends Bone by further dividing both the trainable matrix and $W_{0}$ into smaller shards, where each shard interacts with the trainable matrix $D$ through a linear transformation, achieving nonlinear independence among shard updates.
-The specific formula is as follows:
-$W_{k/r,d/r,r,r}$ $=Reshape(W_{d,k})$ (4)
-$bone_{d/r,r,r}$ $=Reshape(bone_{r,d})$
-$\Delta W_{d,k}$ $=Reshape(W_{k/r,d/r,r,r}bone_{d/r,r,r}+bone_{d/r,r,r})$
-It is evident that Bat can be defined in many other ways, as illustrated in the figure 4.
-In Section 6, we compared the performance of different partitioning strategies for the structure.
-The formula for Bone-row is as follows:
-$W_{d/r,k/r,r,r}$ $=Reshape(W_{d,k})$ (5)
-$bone_{k/r,r,r}$ $=Reshape(bone_{r,k})$
-$\Delta W_{d,k}$ $=Reshape(W_{d/r,k/r,r,r}bone_{k/r,r,r}+bone_{k/r,r,r})$
-Figure 4: Diagram of Bat Structures in Different Design Spaces.
+Inspired by methods like PiSSA that leverage pre-trained weight matrix information, we propose Block Affine Transformation (Bat) to break update collinearity without adding parameters.
+The key insight is to leverage pre-trained weights $\mathbf{W}_{0}$ as nonlinear projectors:
+1. Tensor Factorization:
+•  Reshape $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$ into 4D tensor $\mathcal{W}_{0}\in\mathbb{R}^{\frac{k}{r}\times\frac{d}{r}\times r\times r}$
+•  Reshape $\mathbf{D}\in\mathbb{R}^{r\times d}$ into $\mathcal{D}\in\mathbb{R}^{\frac{d}{r}\times r\times r}$
+2. Affine Transformation: Compute shard-specific updates via tensor contraction:
+$\Delta\mathcal{W}=\mathcal{W}_{0}\times\mathcal{D}+\mathcal{D}\quad\in\mathbb{R}^{\frac{k}{r}\times\frac{d}{r}\times r\times r}$
+where ×$\times_{3}$ denotes contraction along the third dimension.
+3. Reconstruction: Reshape $\Delta\mathcal{W}$ to obtain full update matrix:
+$\Delta\mathbf{W}=\operatorname{Reshape}(\Delta\mathcal{W})\in\mathbb{R}^{d\times k}$
+Bat allows for flexible configuration of different dimensional transformation strategies based on the settings of $\mathbf{D}$. For example:
+Bat-Row: Reshape $\mathbf{W}_{0}$ into $\mathcal{W}\in\mathbb{R}^{\frac{d}{r}\times\frac{k}{r}\times r\times r}$ and $\mathbf{D}\in\mathbb{R}^{r\times k}$ into $\mathcal{D}\in\mathbb{R}^{\frac{k}{r}\times r\times r}$
+Bat-Col: Reshape $\mathbf{W}_{0}$ into $\mathcal{W}\in\mathbb{R}^{\frac{k}{r}\times\frac{d}{r}\times r\times r}$ and $\mathbf{D}\in\mathbb{R}^{r\times d}$ into $\mathcal{D}\in\mathbb{R}^{\frac{d}{r}\times r\times r}$
+The term $\mathcal{W}_{0}\times\mathcal{D}$ introduces shard-dependent perturbations proportional to $\mathbf{W}_{0}$’s singular vectors, breaking the collinearity enforced by Bone’s shared $\mathbf{D}$.
+As shown in Table 7, Bat enhances the model’s performance.
 4 Experiments
@@
 The experiments were conducted on 4×NVIDIA 4090 24G GPUs.
-4.1 Experiments on Natural Language Understanding
-Models and Datasets
-We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC.
-Performance is evaluated on the development set using accuracy as the primary metric.
-Implementation Details
-The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU.
-Each experiment is conducted with 3 different random seeds, and the average performance is reported.
-Results
-As shown in Table 1, Bone demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and Pissa.
 Table 1: The results of fine-tuning RoBERTa-base using Bone and various LoRA variants were compared on a subset of the GLUE benchmark.
@@
 Bone 0.236% $85.71_{\pm 0.32}$ $93.60_{\pm 0.07}$ $\mathbf{72.86_{\pm 3.13}}$ $\mathbf{91.43_{\pm 0.76}}$ $\mathbf{88.14_{\pm 0.60}}$
-4.2 Experiment on Natural Language Generation
-Models and Datasets
-To verify the generalizability of Bone, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math, code, and chat.
-1.
-Math: We trained our model on a 395k subset of MetaMathQA Yu et al.
-(2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K Cobbe et al.
-(2021) and MATH  Yu et al.
-(2023), with higher complexity and diversity.
-2.
-Code: We train our model on a 100k subset of CodeFeedback Zheng et al.
-(2024b), a high-quality code instruction dataset, removing explanations after code blocks.
-The model is tested on HumanEval  Chen et al.
-(2021).
-3.
-Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct Xu et al.
-(2024).
-We test our model on the MT-Bench dataset  Zheng et al.
-(2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects.
-We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
-Figure 5: The image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation.
-Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings.
-Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
-Implementation Details
-The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
-It is worth noting that the weights of LLaMA2-7B are not fully symmetric, making it impossible to perfectly align the trainable parameters when comparing Bone and LoRA.
-To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Bone to 64, ensuring that Bone uses fewer parameters than LoRA to demonstrate its superiority.
-Each experiment is conducted with 2 different random seeds, and the average performance is reported.
-Result
-The results, as shown in Table  2 and Figure  5, demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
-Bone demonstrates outstanding performance across three different tasks.
-On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants.
-On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
 Table 2: We fine-tuned LLMs using Bone and various LoRA variants, and evaluated performance on GSM8k, Math, HumanEval, and MT-Bench.
@@
 Bone 88.1M 66.94 18.85 41.76 -
+4.1 Experiments on Natural Language Understanding
+Models and Datasets
+We fine-tune the RoBERTa-base model on several datasets from the GLUE benchmark, including MNLI, SST-2, CoLA, QNLI, and MRPC.
+Performance is evaluated on the development set using accuracy as the primary metric.
+Implementation Details
+The experimental hyperparameter settings were aligned with those in the LoRA repository, but training was conducted using a single 4090 GPU.
+Each experiment is conducted with 3 different random seeds, and the average performance is reported.
+Results
+As shown in Table 1, Bone demonstrates outstanding performance, particularly on the CoLA dataset, where it exhibits significantly faster convergence and superior data-fitting capabilities, far surpassing LoRA and Pissa.
+4.2 Experiment on Natural Language Generation
+Models and Datasets
+To verify the generalizability of Bone, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math, code, and chat.
+1.
+Math: We trained our model on a 395k subset of MetaMathQA (Yu et al., 2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (Cobbe et al., 2021) and MATH  (Yu et al., 2023), with higher complexity and diversity.
+2.
+Code: We train our model on a 100k subset of CodeFeedback (Zheng et al., 2024b), a high-quality code instruction dataset, removing explanations after code blocks.
+The model is tested on HumanEval  (Chen et al., 2021).
+3.
+Chat: We train our model on a 70k subset of WizardLM-Evol-Instruct (Xu et al., 2024).
+We test our model on the MT-Bench dataset  (Zheng et al., 2024a), which consists of 80 multi-turn questions designed to assess LLMs on multiple aspects.
+We used GPT-4o to judge the quality of responses, as shown in lm-sys/FastChat.
+Implementation Details
+The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used.
+It is worth noting that the weights of LLaMA2-7B are not fully symmetric, making it impossible to perfectly align the trainable parameters when comparing Bone and LoRA.
+To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of Bone to 64, ensuring that Bone uses fewer parameters than LoRA to demonstrate its superiority.
+Each experiment is conducted with 2 different random seeds, and the average performance is reported.
+Result
+The results, as shown in Table  2 and Figure  5, demonstrate that Bone outperforms other PEFT methods in terms of convergence speed, data fitting, and generalization capabilities.
+Bone demonstrates outstanding performance across three different tasks.
+On LLaMA2-7B, Bone achieves results that surpass Pissa, despite using fewer parameters than LoRA and its variants.
+On RWKV6-7B, Bone and LoRA have the same number of trainable parameters, yet Bone consistently outperforms LoRA and its variants across all tasks.
+Figure 5: The image shows the loss curve for LLaMA2-7B fine-tuned on the MetaMathQA dataset, with the first 100 steps highlighted for closer observation.
+Comparing the loss curves reveals that Bone demonstrates superior fitting ability across various architectures and parameter settings.
+Additionally, Bone exhibits a rapid decrease in loss within the first 100 steps, highlighting its effectiveness.
 4.3 Effect of Rank $r$
@@
 4.4 Bone vs Bat
+In this section, we have verified the effectiveness of non-linear updates by comparing the performance of Bone and Bat, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K.
+As shown in the table, with the same number of trainable parameters, the metrics of non-linear updates are better, successfully breaking the original limits.
 Table 4: Comparing Bone, Bat on math tasks
@@
 Bat 55.1M 42.76 6.60
-In this section, we have verified the effectiveness of non-linear updates by comparing the performance of Bone and Bat, we fine-tuned LLaMA2-7B and RWKV6-7B using both Bone and Bat on the MetaMathQA dataset and evaluated their performance on Math and GSM8K.
-As shown in the table, with the same number of trainable parameters, the metrics of non-linear updates are better, successfully breaking the original limits.
 4.5 Ablation experiments on different grouping methods for Bat
 In this subsection, we explore the impact of different grouping methods in the Bat structure on model fine-tuning performance.
-Due to structural differences in the weight matrix W, the Bat-free grouping requires manual configuration, which is inconvenient.
+Due to structural differences in the weight matrix $\mathbf{W}$, the Bat-free grouping requires manual configuration, which is inconvenient.
 Therefore, this subsection only compares row-wise and column-wise grouping, both of which can be easily extended to any structure.
@@
 Bat-col 55.1M 25.25 3.09
+5 Resource and efficiency
 The experimental results, shown in Table  5, indicate that the difference between the two is minimal, with both performing well.
@@
 However, this doesn’t imply that different grouping methods are meaningless.
-As more LLMs begin to use techniques like GQA and MLA Liu et al.
-(2024a) to reduce KV cache overhead Dai et al.
-(2024); Lee et al.
-(2024); Shazeer (2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
-5 Resource and efficiency
+As more LLMs begin to use techniques like GQA and MLA (Liu et al., 2024a) to reduce KV cache overhead (Dai et al., 2024; Lee et al., 2024; Shazeer, 2019), the main weight matrices become smaller, and Bat will need to adjust its grouping or employ other techniques to adapt to these new technologies.
 Table  6 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA, Bone, and Bat on a single 4090 GPU.
@@
 Table 6: Resource and efficiency
-Model Strategy Trainable GPU Memory Token throughput
-RWKV-3B LoRA 55.1M 12074 MB 3.62 kt/s
-Bone 55.1M $\mathbf{11052}$ MB $\mathbf{3.99}$ kt/s
+devices Strategy Trainable Memory throughput
+1*gpu LoRA 55.1M 12074 MB 3.62 kt/s
+Bone 55.1M 11052 MB 3.99 kt/s
 Bat 55.1M 22978 MB 2.16 kt/s
-RWKV-3B LoRA(use checkpoint) 55.1M 4*15328 MB 15.6 kt/s
-Bone(use checkpoint) 55.1M $\mathbf{4*15304}$ MB $\mathbf{16.0}$ kt/s
-Bat(use checkpoint) 55.1M 4*15305 MB 14.2 kt/s
+4*gpu LoRA 55.1M 4*15328 MB 15.6 kt/s
+Bone 55.1M 4*15304 MB 16.0 kt/s
+Bat 55.1M 4*15305 MB 14.2 kt/s
 6 Conclusion
@@
 This approach significantly reduces resource overhead and opens up new directions for PEFT techniques.
-Building on the DiSHA framework, we introduced Bone, a broad and efficient method comparable to LoRA.
+To enhance the adaptability and efficiency of DiSHA, we proposed Block Affine Efficient Computation (Bone).
 Extensive experiments demonstrated the superiority of Bone, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
-To address the potential issue of linear correlation in shard updates within Bone and the DiSHA framework, we proposed Bat.
-By applying a linear transformation between the trainable matrix and the pre-trained weights, Bat ensures nonlinear independence in the updates of different shards, successfully breaking through the structural limitations of Bone
+Bat breakthrough to overcome the collinear update bottleneck inherent in both DiSHA and Bone.
+By integrating pre-trained weight information through linear transformation.
+Bat induces nonlinearly independent shard updates without extra parameters.
+Collectively, our framework redefines the design principles of efficient adaptation: DiSHA provides theoretical grounding for dimension-wise decomposition, Bone delivers practical efficiency parity with LoRA, and Bat unlocks higher-order update diversity previously unattainable in low-rank methods.
+This progression demonstrates that parameter efficiency need not come at the cost of expressivity, paving the way for adaptive fine-tuning in resource-constrained environments.
 7 Future work
@@
 Finally, I sincerely appreciate the meticulous guidance provided by the anonymous online mentor, Professor Cormen, whose help has been indispensable.
+Appendix A Appendix
+A.1 Gradient Norm
+To investigate the factors affecting the convergence speed of LLM training, we conducted extensive experiments.
+In the gradient norm analysis, we observed that a larger initial gradient norm leads to faster loss convergence.
+Initially, we compared full fine-tuning (FT), LoRA, and PiSSA, as shown in the figure 1(The figure illustrates the initial gradient norm of model training for different structures at various ranks.
+Notably, FT (full fine-tuning), Frozen FT, and BigLoRA remain unaffected by rank, as their trainable parameters remain constant.) and figure 5.
+Many studies suggest that LoRA’s slow convergence speed is likely due to the small gradients caused by the zero initialization of either multiplicand matrix.
+However, we hypothesize that LoRA’s slow convergence could also be attributed to structural issues.
+To test this hypothesis, we introduced two additional experiments: Frozen FT and BigLoRA.
+Both experiments freeze the pre-trained weights $\mathbf{W}_{0}$, but they differ in design.
+Frozen FT replaces $\mathbf{W}_{0}$ with a trainable matrix $\mathbf{W}_{t}$ of the same dimensions, initialized to zero.
+BigLoRA, on the other hand, sets the rank to 1024, ensuring that its trainable parameter count exceeds both Frozen FT and FT.
+We observed an intriguing phenomenon: although the initial gradient norm of LoRA and PiSSA increases with larger ranks, BigLoRA, despite having significantly more trainable parameters than Frozen FT or FT, still only has an initial gradient norm of 12.64.
+Meanwhile, Frozen FT, despite being initialized to zero, has an initial gradient norm only slightly smaller than FT.
+This suggests that there may be simpler structural solutions that can naturally unlock lower intrinsic ranks and achieve faster convergence without the need for additional complexity.
+Finally, we added the initial gradient data of Bone to represent the performance of the DiSHA framework.
+A.2 Efficiency Analysis
+Table 7: Comparing Bone, Bat on math tasks
+Operation Memory Complexity FLOPs
+LoRA ($\mathbf{BA}x$) $O(dr+rk)$ $O(blr(d+k))$
+DiSHA ($\operatorname{expand}(\mathbf{D})x$) $O(dk)$ $O(bldk)$
+Bone ($\mathbf{D}^{\top}\mathbf{S}$) $O(dr)$ $O(blr(d+\frac{k}{r}))$
+Memory Reduction: Bone saves $\frac{k}{r}$ memory over DiSHA when $r\ll k$.
+FLOPs Advantage: Bone’s blockwise summation replaces $O(bldk)$ dense multiplication with $O(blk)$ summation and $O(bldr)$ projection, achieving up to 2.0× speedup (see Table 6).